### v2

# CoCoA: A Generalized Approach to Uncertainty Quantification by Integrating Confidence and Consistency of LLM Outputs

Abstract
Uncertainty quantification (UQ) methods for Large Language Models (LLMs) encompasses a variety of approaches, with two major types being particularly prominent: information-based, which focus on model confidence expressed as token probabilities, and consistency-based, which assess the semantic relationship between multiple outputs generated using repeated sampling. Several recent methods have combined these two approaches and shown impressive performance in various applications. However, they sometimes fail to outperform much simpler baseline methods. Our investigation reveals distinctive characteristics of LLMs as probabilistic models, which help to explain why these UQ methods underperform in certain tasks. Based on these findings, we propose a new way of synthesizing model confidence and output consistency that leads to a family of efficient and robust UQ methods. We evaluate our approach across a variety of tasks such as question answering, abstractive summarization, and machine translation, demonstrating sizable improvements over state-of-the-art UQ approaches.
1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing (NLP), enabling advancements in information retrieval, question answering, machine translation, and other language-driven applications. As these models become an integral part of everyday life, ensuring the reliability of their outputs is crucial, especially in high-stakes scenarios where errors or uncertainty can have serious consequences. One way to address this challenge is through uncertainty quantification (UQ), which measures how confident a model is in its outputs, and makes possible the rejection of generations with a high risk of being incorrect.
UQ for LLMs is a rapidly advancing research area, with new methods for estimating uncertainty emerging each year. The large portion of novel techniques is based on two fundamental approaches: information-theoretic analysis and the assessment of output consistency.
Figure 1: Example of inconsistent probabilities assigned to semantically identical answers by an LLM, demonstrating the limitation of relying solely on sequence-level information.
Information-theoretic methods quantify the confidence of a model by analyzing the probability distributions it induces for predictions (Malinin and Gales, 2021; Fomicheva et al., 2020). A key limitation of these methods is that they cannot account for the semantic variability across multiple possible outputs for the same input. Specifically, the model may generate answers with the same meaning but with very different assigned probabilities; see Figure 1. LLMs are trained to predict the next token in a sequence based on patterns observed in vast amounts of data, resulting in varying probabilities for semantically equivalent output sequences.
In contrast, consistency-based methods directly analyze the semantic relationships between the sampled outputs (Lin et al., 2024; Fomicheva et al., 2020). Information-theoretic and consistency-based methods have complementary strengths: the former provides insights into the model’s internal confidence, while the latter captures the uncertainty as objective variability of meaning among sampled outputs. For this reason, recent state-of-the-art methods aimed to unify these approaches (Kuhn et al., 2023; Duan et al., 2024). Although such methods show good performance in various applications, they sometimes fail to outperform their simpler counterparts in certain scenarios (Vashurin et al., 2024).
Our investigation revealed distinctive characteristics of LLMs as probabilistic models, shedding light on why current state-of-the-art UQ methods that attempt to integrate both approaches often underperform in certain tasks. Specifically, we highlight the complexity of the token prediction process and the absence of a unified framework that simultaneously addresses model confidence and output variability, both of which can limit the effectiveness of existing UQ techniques. This insight drives our proposal for a novel family of methods that integrate model confidence with output consistency, resulting in more efficient and robust UQ techniques. Our approach combines the strengths of both information-based and consistency-based methods, providing a more comprehensive and accurate assessment of uncertainty.
Our main contributions can be summarized as follows.
  • We identify key limitations in current UQ methods for LLMs, particularly in addressing both token- and sequence-level confidence and output consistency.
  • We present a family of Confidence and Consistency-based Approaches (CoCoA) to UQ, offering a new way to merge information- and consistency-based measures for uncertainty quantification in LLMs.
  • We evaluate our approaches across a variety of NLP tasks, including question answering, summarization, and translation. Our experiments demonstrate sizable improvements in the reliability and robustness of UQ compared to state-of-the-art methods.
Figure 2: Illustration of the method: the LLM generates a response, evaluates the similarity to alternatives, computes the uncertainty, and combines the uncertainty score with the similarity measures. High similarity to alternatives reduces the uncertainty, while low similarity keeps it high.
2 Background
In this section, we introduce key concepts related to uncertainty quantification for LLMs, outline existing methods, discuss their limitations, and highlight the motivation for our approach.
First and foremost, it is important to establish the concept of an uncertainty function. Let $\mathbf{y}=f(\mathbf{x})$ denote the output of an LLM given an input sequence $\mathbf{x}$. The model defines a probabilistic output distribution $p(\mathbf{y}\mid\mathbf{x})$, from which outputs can be sampled. An uncertainty function $U$ is a mapping that quantifies the level of uncertainty $u$ associated with the output of a model $\mathbf{y}$, conditioned on the input sequence $\mathbf{x}$, which we denote as
$u=U(\mathbf{y}\mid\mathbf{x}).$ (1)
With this notion in mind, we now discuss existing methods for uncertainty quantification, outline their limitations, and describe how our approach addresses them.
2.1 Single-Sequence Information-Based Methods
Information-based methods rely on a single sample from the LLM and estimate the uncertainty of the generated sequence by aggregating the uncertainty scores of individual tokens. One of the simplest techniques of this kind is Maximum Sequence Probability (MSP):
$u_{\text{MSP}}=-\log p\bigl{(}\mathbf{y}\mid\mathbf{x}\bigr{)}.$ (2)
Several other measures fall into this category, including Perplexity and Mean Token Entropy (Fomicheva et al., 2020); see Appendix C.1 for details. While using only a single sample makes them computationally efficient, these techniques face three major challenges:
  1. 1.
    First, LLMs provide us with the probability of a specific answer, even though the same meaning could often be expressed in multiple ways. Therefore, to obtain a proper probability for the meaning of an answer, we need to marginalize over its various possible rephrasings. However, this is not feasible if we generate only a single sample.
  2. 2.
    Second, answers of LLMs in general have very low probabilities due to the size of the vocabulary and the length of the generated sequences. Therefore, their estimates by LLM are not very reliable in general.
  3. 3.
    Third, these methods do not provide information about the flatness or the variability of the answer distribution.
2.2 Consistency-Based Methods
The aforementioned issues lead to the development of consistency-based methods based on repetitive sampling from the LLM. Consider that we have sampled a set of outputs $\bigl{\{}\mathbf{y}^{(i)}\bigr{\}}_{i=1}^{M}$, where ${\mathbf{y}^{(i)}\sim p(\mathbf{y}\mid\mathbf{x})}$. Consistency-based uncertainty quantification methods rely only on the diversity of answers $\mathbf{y}^{(i)}$ sampled from the LLM. The idea is that if the model outputs similar answers for the same prompt over and over again, it is confident in its predictions; otherwise, it is uncertain. These techniques do not require knowledge about the probability distribution of the tokens and can be applied in the black-box setting, when only the generated tokens are available. This case is quite common when LLMs are deployed as a service and are accessible through a limited API.
Formally, given $M$ samples from the model, consistency-based methods compute a similarity matrix $G$, where each element $g_{ij}$ represents some form of similarity between the sampled outputs $\mathbf{y}^{(i)}$ and $\mathbf{y}^{(j)}$:
$g\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}\in[0,1].$ (3)
The value $g\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}=1$ indicates the complete equivalence between $\mathbf{y}^{(i)}$ and $\mathbf{y}^{(j)}$, and $g\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}=0$ indicates that there is no similarity.
Similarity could be computed in various ways. For example, Lexical Similarity (Fomicheva et al., 2020) is the surface form similarity, which calculates the overlap of words or phrases in the generations. More advanced techniques propose various methods for taking into account the semantic similarity of the generated answers by hard or soft clustering (Lin et al., 2024). For example, the Degree Matrix approach considers a similarity matrix $G$, which is computed using a model for Natural Language Inference (NLI), which predicts the probabilities of entailment, ${p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})$, and contradiction, ${p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$, between pairs of sentences $\mathbf{y}$ and $\mathbf{y}^{\prime}$. The similarity between two sequences is then defined as either $g_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})={p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})\quad\text{or}\quad g_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})=1-{p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$. Lin et al. (2024) proposes to compute an averaged similarity matrix as $g_{ij}=\bigl{(}g\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}+g\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(i)}\bigr{)}\bigr{)}/2$. The diagonal matrix $D$ is then defined with elements $D_{ii}=\sum_{j=1}^{M}g_{ij}$. The uncertainty from this soft clustering of meanings can then be defined as follows:
$U_{\mathrm{DegMat}}=1-\frac{\mathrm{trace}(D)}{M^{2}}.$ (4)
The advantage of these techniques is that by generating multiple samples and analyzing their semantic similarity, they can obtain empirical probabilities for meanings instead of individual answers. The main drawback is that they discard the useful information that comes from the probability distribution represented by the LLM, including estimates of probabilities of specific answers.
2.3 Information-Based Methods with Repeated Sampling
The natural idea is to somehow benefit from having multiple samples from the model while using important information contained in the output probabilities estimated by an LLM. Below, we examine several approaches that have sought to achieve this.
Averaging uncertainties.
The uncertainty scores can be aggregated using simple Monte Carlo averaging:
$u_{\text{MC}}=\frac{1}{M}\sum_{i=1}^{M}u_{i}.$ (5)
For the case when using the MSP uncertainty measure, i.e., when $u_{i}=-\log p\bigl{(}\mathbf{y}^{(i)}\mid\mathbf{x}\bigr{)}$, we obtain $u_{\text{MC}}=-\frac{1}{M}\sum_{i=1}^{M}\log p\bigl{(}\mathbf{y}^{(i)}\mid\mathbf{x}\bigr{)}$. The other notable example is the Monte Carlo Sequence Entropy (Kuhn et al., 2023).
While simple averaging represents a natural way to aggregate uncertainties, it has certain issues related to the nature of LLMs. First of all, in the vast majority of applications, an LLM-based system should produce a single output $\mathbf{y}_{*}$ for an input query. When we consider $u_{\text{MC}}$, we essentially perform averaging of uncertainties of different sequences, thus somewhat assessing the uncertainty related to the entire generative distribution $p(\mathbf{y}\mid\mathbf{x})$ for the input $\mathbf{x}$, but not for a particular generated sequence $\mathbf{y}_{*}$. This averaged uncertainty might not be adequate for this particular sequence and, remarkably, often performs worse than the uncertainty $u_{*}=U(\mathbf{y}_{*}\mid\mathbf{x})$, which is related solely to the output $\mathbf{y}_{*}$. Moreover, although intuitive, this naïve aggregation method assumes that all outputs contribute equally to the final uncertainty estimate, regardless of their semantic relationships. This can lead to inconsistencies when semantically equivalent outputs have varying uncertainty scores or when outputs with low similarity are treated as equally important.
Semantically weighted averaging.
The basic idea of aggregation approaches like Semantic Entropy (Kuhn et al., 2023) or SAR (Duan et al., 2024) is to perform a weighted averaging of output probabilities and give more weight to sequences semantically similar to the response shown to a user. All recently proposed techniques, such as SAR and Semantic Entropy, can be unified into a semantically-aware Generalized Monte Carlo uncertainty estimate, defined as
$u_{\text{GMCU}}=\frac{1}{M}\sum_{i=1}^{M}h\Biggl{(}\sum_{j=1}^{M}g_{ij}\,p_{j}\Biggr{)}.$ (6)
Here, the inner summation aggregates sequence probabilities $p_{j}$ weighted by their semantic similarity to the $i$-th output, and the outer summation averages these contributions across all samples. The function $h(\cdot)$ provides an additional layer of flexibility, transforming the reweighted uncertainty scores, making the method a generalized framework for uncertainty quantification. Existing methods, such as Semantic Entropy and SAR, can be considered as special cases of this more comprehensive approach, where the functions $h$ and $g$ are chosen appropriately.
Unfortunately, methods that fall under GMCU, while offering benefits, also inherit the aforementioned issues from both categories of methods:
  1. 1.
    The term $\sum_{j=1}^{M}g_{ij}\,p_{j}$ aims to average the probabilities of semantically similar sequences to obtain more robust estimate of the probability. However, due to the extreme instability of the LLM probabilities, as shown in Figure 1, the aggregated probabilities often perform worse than non-aggregated baselines.
  2. 2.
    The outer summation in (6), similarly to the case of simple Monte Carlo averaging (5), often fails to outperform the uncertainty $u_{*}=U(\mathbf{y}_{*}\mid\mathbf{x})$ of a single generated sequence $\mathbf{y}_{*}$.
Let us note that all uncertainty functions discussed so far have the following properties:
  • Non-Negativity. The uncertainty function produces nonnegative values, that is, $U(\mathbf{y})\geq 0$ for all $\mathbf{y}$.
  • Monotonicity. Higher values of the uncertainty function $U$ indicate higher uncertainty, i.e., if output $\mathbf{y}^{(1)}$ is considered more uncertain than output $\mathbf{y}^{(2)}$, then $U\bigl{(}\mathbf{y}^{(1)}\bigr{)}\geq U\bigl{(}\mathbf{y}^{(2)}\bigr{)}$.
These properties become important to perform the kind of synthesis of confidence and consistency we propose in the following section.
3 CoCoA: Bridging Confidence and Consistency for Better Uncertainty Quantification
We start by summarizing the benefits and drawbacks of various uncertainty quantification approaches discussed above:
  1. 1.
    Both information-based and (semantic) consistency-based methods provide grounded and useful uncertainty quantification measures.
  2. 2.
    Output probabilities $p(\mathbf{y}^{(j)}\mid\mathbf{x}),\,j=1,\dots,M$ might have substantially different values for semantically equivalent outputs, which questions the usefulness of (weighted) averaging these probabilities for uncertainty quantification.
  3. 3.
    For various methods based on the aggregation over multiple samples, the result might be suboptimal due to the noise related to the averaging over all generated outputs. Focusing solely on a particular output sequence and its relation to other generated outputs might be beneficial.
In what follows, we present a family of UQ Confidence and Consistency-based Approaches (CoCoA), offering a new way to merge information- and consistency-based measures for uncertainty quantification in LLMs.
Let us consider an actual output sequence $\mathbf{y}_{*}$ and a set of sampled sequences $\mathbf{y}^{(i)},\,i=1,\dots,M$. Here, $\mathbf{y}_{*}$ might be one of the sequences $\mathbf{y}^{(i)}$ or might be generated separately. In what follows, we will consider several possible cases, including $\mathbf{y}_{*}$ being a random sequence from a set $\{\mathbf{y}^{(i)}\}$, $\mathbf{y}_{*}$ being a sequence from a set $\{\mathbf{y}^{(i)}\}$ having the highest probability, and, finally, $\mathbf{y}_{*}$ being a sequence found via the beam search procedure.
First, consider an information-based uncertainty score of the output $\mathbf{y}_{*}$:
$u_{*}^{\text{info}}=U^{\text{info}}(\mathbf{y}_{*}\mid\mathbf{x}),$ (7)
where $U^{\text{info}}$ might be MSP, perplexity, mean token entropy, or another uncertainty measure related solely to the generated sequence $\mathbf{y}_{*}$.
We quantify the consistency-based uncertainty via a direct measurement of the semantic similarity of generated sequence $\mathbf{y}_{*}$ to sampled sequences:
$u_{*}^{\text{cons}}=\frac{1}{M}\sum_{i=1}^{M}(1-g_{*i}),$ (8)
where $g_{*i}=g\bigl{(}\mathbf{y}_{*},\mathbf{y}^{(i)}\bigr{)}$. This formulation satisfies the desired properties of the uncertainty function – that is their values are nonnegative and their values increase with increased inconsistency (decreasing value of $g_{*i}$). In our ablation study, we will show that such an uncertainty measure reliably outperforms consistency-based measures that aggregate the pairwise similarities of all the samples (see Appendix B.3).
Finally, we need to aggregate $u_{*}^{\text{info}}$ and $u_{*}^{\text{cons}}$ into a single uncertainty measure. We propose to aggregate them in a multiplicative way:
$u_{*}^{\texttt{CoCoA}}=u_{*}^{\text{info}}\cdot u_{*}^{\text{cons}}$ (9)
This formulation preserves the non-negativity and the monotonicity properties while integrating both global (semantic) and local (model-specific) uncertainty signals. It ensures that uncertainty is amplified for sequences that are both intrinsically uncertain (high $u_{*}^{\text{info}}$) and semantically inconsistent with the dataset (high $u_{*}^{\text{cons}}$), while keeping it low for the opposite scenario (see Figure 2).
Although the choice of the multiplicative aggregation function CoCoA is heuristic, it provides a practical and effective way to combine information- and consistency-based uncertainty signals in LLMs. In our ablation study, we also compare the multiplicative formulation in (9) to a simpler additive variant, $u_{*}^{\text{info}}+u_{*}^{\text{cons}}$ (see Appendix B.3). Empirically, the multiplicative combination is better at capturing the joint impact of both information-based and consistency-based uncertainty, yielding more reliable estimation across all tasks.
4 Experiments
4.1 Experimental Setup
To evaluate the effectiveness of the proposed method, we extended the LM-Polygraph library (Vashurin et al., 2024; Fadeeva et al., 2023) by implementing our approach within its framework. Since the library already includes tools for calculating other uncertainty scores, it provided a convenient and efficient environment for setting up and running experiments. The primary objective of our experiments is to evaluate whether our method offers improved performance in key tasks such as question answering (QA), summarization (SUM), and machine translation (MT), compared to existing baselines.
Figure 3: Prediction-Rejection Ratio (PRR) Curve. The curve illustrates the quality of the non-rejected predictions as a function of the rejection rate. Oracle represents the optimal rejection strategy, Random is a random rejection, and UE is rejection based on the evaluated uncertainty estimation method.
Datasets.
For QA, we selected diverse datasets to capture a variety of challenges: TriviaQA (Joshi et al., 2017), an open-domain factual QA dataset; CoQA (Reddy et al., 2019), a conversational QA benchmark requiring multi-turn contextual understanding; MMLU (Hendrycks et al., 2021), a multi-task dataset spanning 57 topics to test broad knowledge; and GSM8k (Cobbe et al., 2021), which focuses on grade-school math problems requiring logical reasoning. For translation, we evaluated our method on WMT14 French-English (Bojar et al., 2014) and WMT19 German-English (Barrault et al., 2019). Finally, for summarization, we used XSum (Narayan et al., 2018), a dataset of complex documents paired with concise abstractive summaries.
For all datasets, we follow (Vashurin et al., 2024) for selecting the subsets, for prompt formatting, and for number and sourcing of few-shot examples.
Models.
We evaluated our method using the base versions of three open-weights language models: LLaMA 3.1 8B (Touvron et al., 2023), Mistral 7B (Jiang et al., 2023), and Falcon 3 7B (Team, 2024). The open-source nature of these models provides full access to their token probabilities, which are essential for implementing our UQ method. For all models, we consider base versions, without instruction-tuning.
Metric Llama Mistral Falcon
QA NMT SUM QA NMT SUM QA NMT SUM
MCSE 0.357 0.380 0.192 0.453 0.406 0.162 0.460 0.409 0.128
MCNSE 0.380 0.429 0.186 0.466 0.489 0.196 0.530 0.424 0.153
Semantic Entropy 0.397 0.411 0.194 0.482 0.438 0.164 0.479 0.440 0.134
SAR 0.479 0.506 0.159 0.542 0.576 0.175 0.590 0.488 0.193
DegMat 0.422 0.342 0.191 0.465 0.425 0.205 0.543 0.386 0.177
EigValLaplacian 0.388 0.274 0.190 0.426 0.366 0.197 0.498 0.336 0.174
MSP 0.395 0.376 0.464 0.444 0.252 0.330 0.343 0.381 0.099
$\text{CoCoA}_{MSP}$ 0.484 $\uparrow$ 0.607 $\uparrow$ 0.484 $\uparrow$ 0.526 $\uparrow$ 0.721 $\uparrow$ 0.366 $\uparrow$ 0.529 $\uparrow$ 0.631 $\uparrow$ 0.210 $\uparrow$
PPL 0.532 0.563 0.458 0.587 0.686 0.365 0.627 0.589 0.275
$\text{CoCoA}_{PPL}$ 0.571 $\uparrow$ 0.617 $\uparrow$ 0.450 0.613 $\uparrow$ 0.745 $\uparrow$ 0.372 $\uparrow$ 0.647 $\uparrow$ 0.648 $\uparrow$ 0.310 $\uparrow$
NMTE 0.477 0.469 0.449 0.559 0.637 0.350 0.602 0.492 0.186
$\text{CoCoA}_{NMTE}$ 0.548 $\uparrow$ 0.579 $\uparrow$ 0.451 $\uparrow$ 0.600 $\uparrow$ 0.720 $\uparrow$ 0.373 $\uparrow$ 0.641 $\uparrow$ 0.614 $\uparrow$ 0.289 $\uparrow$
Table 1: Results for Evaluated Sequence – Best Sample: Mean PRR across datasets for each task. The best performing method is in bold, and the second-best is underscored. Arrows indicate improvement in CoCoA over the base version.
Similarity Function.
To measure the similarity, we use the RoBERTa-large cross-encoder model, fine-tuned on the Semantic Textual Similarity benchmark dataset (Liu et al., 2019; Reimers and Gurevych, 2019; Cer et al., 2017). This model is widely regarded as one of the most reliable and commonly used approaches for evaluating sentence similarity. The cross-encoder processes two sequences jointly and directly outputs a similarity score ranging from 0 to 1, providing a nuanced measure.
Appendix B.2 contains comparative experiments with cross-encoder and other choices of the similarity function, substantiating this choice.
Baselines.
We compare the performance of the proposed method against a diverse set of baseline and state-of-the-art UQ scores, including confidence-based, consistency-based, and hybrid approaches. Information-based methods include Maximum Sequence Probability (MSP), Perplexity (PPL), Mean Token Entropy (MTE), Monte Carlo Sequence Entropy (MCSE), and Monte Carlo Normalized Sequence Entropy (MCNSE). Consistency-based methods include the Degree Matrix (DegMat) and the Sum of Eigenvalues of the Graph Laplacian (EigValLaplacian). Finally, Hybrid methods include Semantic Entropy and SAR. All formulations for these baselines can be found in Appendix C.
Evaluation measure.
As our evaluation measure, we choose the Prediction Rejection Ratio (PRR), which measures the effectiveness of the uncertainty scores for identifying high-quality predictions (Malinin and Gales, 2021). PRR operates by progressively rejecting predictions with uncertainty scores above a threshold $a$ and observing how the average quality $Q(f(\mathbf{x}_{i}),\mathbf{y}_{i})$ of the remaining predictions changes. The metric is calculated as the ratio of two areas: the area between the Prediction Rejection (PR) curves for the evaluated uncertainty score and a random baseline, and the area between the oracle (the ideal uncertainty score that perfectly ranks instances by quality) and the random baseline. Formally, PRR is defined as follows:
$PRR=\frac{\text{AUC}_{\text{unc}}-\text{AUC}_{\text{rnd}}}{\text{AUC}_{\text{oracle}}-\text{AUC}_{\text{rnd}}}.$ (10)
Higher PRR values indicate better alignment of uncertainty scores with prediction quality, approaching the performance of an oracle. To ensure practical applicability, we compute PRR only up to a rejection threshold of 50%, preventing cases where excessive rejection artificially inflates quality measures. Figure 3 gives a visual representation of the PRR calculation, highlighting the relationship between the uncertainty threshold and the quality measures.
Metric Llama Mistral Falcon
QA NMT SUM QA NMT SUM QA NMT SUM
MCSE 0.310 0.323 0.033 0.389 0.304 0.007 0.414 0.317 0.159
MCNSE 0.309 0.393 0.022 0.384 0.410 0.009 0.405 0.422 0.108
Semantic Entropy 0.356 0.343 0.033 0.423 0.327 0.008 0.439 0.348 0.164
SAR 0.414 0.455 0.077 0.462 0.435 0.094 0.481 0.458 0.144
DegMat 0.406 0.302 0.081 0.423 0.305 0.137 0.483 0.353 0.201
EigValLaplacian 0.375 0.238 0.079 0.391 0.267 0.132 0.459 0.312 0.201
MSP 0.409 0.399 0.328 0.475 0.383 0.287 0.475 0.356 0.201
$\text{CoCoA}_{MSP}$ 0.451 $\uparrow$ 0.519 $\uparrow$ 0.378 $\uparrow$ 0.509 $\uparrow$ 0.497 $\uparrow$ 0.330 $\uparrow$ 0.511 $\uparrow$ 0.505 $\uparrow$ 0.257 $\uparrow$
PPL 0.381 0.386 0.369 0.424 0.427 0.204 0.456 0.450 0.155
$\text{CoCoA}_{PPL}$ 0.454 $\uparrow$ 0.481 $\uparrow$ 0.387 $\uparrow$ 0.494 $\uparrow$ 0.472 $\uparrow$ 0.286 $\uparrow$ 0.523 $\uparrow$ 0.508 $\uparrow$ 0.229 $\uparrow$
NMTE 0.353 0.382 0.357 0.417 0.438 0.182 0.456 0.473 0.152
$\text{CoCoA}_{NMTE}$ 0.447 $\uparrow$ 0.478 $\uparrow$ 0.380 $\uparrow$ 0.492 $\uparrow$ 0.469 $\uparrow$ 0.288 $\uparrow$ 0.527 $\uparrow$ 0.508 $\uparrow$ 0.228 $\uparrow$
Table 2: Results for Evaluated Sequence – Greedy Sample: Mean PRR across datasets for each task. The best-performing method is shown in bold, and the second-best is underscored. Arrows indicate improvement in CoCoA over the base version.
Quality Measures.
The Predictive Rejection Ratio (PRR) requires an appropriate quality measure for each specific task to effectively evaluate the model output. For question-answering tasks, we use Accuracy to directly evaluate whether the generated answers match the ground truth in short-form QA tasks (e.g., MMLU), and we use the AlignScore between correct answer and generated sequence for assessing the performance for long-form QA tasks (Zha et al., 2023). For summarization tasks, we use AlignScore to measure the alignment between the output summary and the input document. It serves as a quality indicator by evaluating the relevance and the overlap between the generated content with the source text. For translation tasks, we use COMET, as it captures both semantic adequacy and fluency, ensuring that translations are accurate and linguistically appropriate (Rei et al., 2020).
Generation Setup.
We discuss the generation parameters, the decoding strategy and sample selection procedure in depth in Appendix D. In short, we report evaluation results in two distinct setups: greedy decoding and stochastic sampling with focus on the most probable sequence among the generated outputs (best-sample). These two setups offer the highest-quality outputs and are the most reasonable generation approaches in practice.
4.2 Results
Main results.
Tables 1 and 2 show the PRR scores under the best‐sample and greedy generation setup. We calculate a single representative PRR for each task – question answering, neural machine translation (NMT), and summarization (SUM) – by averaging the results across all relevant datasets (e.g., TriviaQA, MMLU, CoQA, GSM8k for QA). This aggregated score provides a concise measure of the performance for each model for each task. Detailed results for each dataset separately can be found in Appendix D.
We can see that our CoCoA methods are the best across all tasks and models in our experiments. They outperform existing consistency‐based and hybrid state‐of‐the‐art approaches, like Semantic Entropy and SAR. In addition, our proposed CoCoA approach consistently surpasses the baseline UE metrics: for example, $\texttt{CoCoA}_{PPL}$ outperforms standard Perplexity, illustrating the advantage of combining token‐level confidence with semantic consistency. This pattern holds for other information‐based metrics as well, demonstrating that using the consistency between multiple sampled outputs reliably enhances uncertainty quantification.
Ablation study.
As a part of our ablation study (see Appendix B.1), we evaluate the performance of the average dissimilarity component ($u_{*}^{\text{cons}}$) independently to assess its effectiveness as a standalone uncertainty measure and to investigate whether it could potentially outweigh the contribution of the information-based component ($u_{*}^{\text{info}}$) in the enriched uncertainty measure. This evaluation enables us to isolate and better understand the complementary roles and the relative importance of each component. Our experiments demonstrate that the combination of consistency- and confidence-based metrics outperforms the pure consistency-based measure on a vast majority of tasks. Notably, in the few cases where the pure consistency messure outperforms the combined approach, the performance difference is minimal. It is possible that a more suitable choice of a similarity measure or confidence-based metric for the task could further improve the performance.
This leads us to the next part of our ablation study, where we investigate the impact of different similarity measures (see Appendix B.2). We find that for some tasks, the similarity score computed by the Cross-encoder does not yield optimal performance. For example, for question-answering tasks on CoQA and Trivia, NLI-derived similarity performs better than the Cross-encoder similarity and outperforms the pure consistency-based uncertainty discussed above.
The next section of our ablation study focuses on alternative forms of combining model confidence $u_{*}^{\text{info}}$ and consistency $u_{*}^{\text{cons}}$ (see Apendix B.3). First, we consider an additive form of combining them: $\mathrm{U}_{\text{AdditiveCoCoA}}=u_{*}^{\text{info}}+u_{*}^{\text{cons}}$. The results show that this additive formulation does not perform as well compared to the multiplicative one. The additive form tends to underemphasize the interaction between the two components, which is critical for capturing the nuanced relationships between confidence and consistency.
We also consider an alternative formulation of the consistency term $u_{*}^{\text{cons}}$, as the average of the full pairwise dissimilarity. In this formulation, $u_{*}^{\text{cons}}$ represents the average inconsistency across all samples rather than focusing solely on the dissimilarity of the evaluated sequence with the other samples. Our experiments demonstrate that this formulation is not very strong. By distributing the consistency computation across all samples, it loses focus on the specific sequence being evaluated.
Lastly, in Appendix B.3, we also consider alternative formulations of the information-based metric that do not rely on logarithmic transformations. While we primarily use logarithms due to their numerical stability, we explore an alternative approach by converting these values back to probabilities and analyzing their impact on uncertainty estimation. Our findings indicate that both formulations exhibit consistent performance and yield similar results. This suggests that while logarithmic transformations enhance numerical stability, the choice between log-based and probability-based formulations does not affect much the overall performance.
5 Limitations
While our proposed CoCoA approach demonstrates robust empirical performance, several important considerations remain.
Task and Domain Dependency.
Our method relies on both an information-based confidence score and a semantic similarity function. The effectiveness of each can vary across models, tasks, and domains. For open-ended tasks with multiple equally valid outputs (e.g., creative generation), consistent rephrasing may inflate the perceived certainty. Conversely, in domains that demand highly precise factual or logical correctness (e.g., math problem solving), small deviations in reasoning can lead to large outcome differences that are not fully captured by a generic similarity measure. Adapting both the confidence measure and the similarity function to specific domains or prompt types is an important direction for future work.
Limited Sample Size.
CoCoA estimates the model’s consistency by sampling multiple outputs and comparing them. In practice, generating a large number of samples can be computationally expensive and may increase the inference latency. Consequently, our experiments (like many sampling-based approaches) rely on relatively small sample sets. Although even a handful of samples can provide a meaningful estimate of consistency, it may not fully capture the diversity of the underlying distribution for certain tasks or for more complex prompts.
Quality Metric.
Finally, the CoCoA’s performance assessment depends on quality metrics (e.g., COMET for machine translation, and Accuracy for QA) that may not capture every nuance of textual outputs. Automatic metrics can have blind spots, particularly in evaluating coherence, factual correctness, or subtle aspects of style. Further refining or extending quality metrics to account for deeper reasoning, factual faithfulness, and stylistic appropriateness would better align uncertainty scores with real-world perceptions of model correctness.
6 Conclusion
We presented CoCoA, a unified approach that integrates Confidence and Consistency for uncertainty quantification in LLMs. By combining token-level confidence scores with semantic similarity between multiple sampled outputs, CoCoA offers a more holistic view of uncertainty than either approach alone. In extensive evaluations on question answering, summarization, and translation, our approach outperformed existing baselines and state-of-the-art UQ methods. Moreover, CoCoA’s flexible design allows easy adaptation to a variety of tasks and settings.
Moving forward, several directions are open for further exploration. These include incorporating more adaptive sampling strategies that efficiently capture the model output space, refining semantic similarity functions for domain-specific tasks, and improving calibration techniques to strengthen the confidence metrics of the model.
Impact Statement
As large language models (LLMs) are increasingly integrated into everyday applications, ensuring the robustness of their outputs is essential, particularly in high-stakes domains such as healthcare and the legal sector. In addition to flagging low-quality responses, clear confidence estimates can help reduce overreliance on automated systems and foster safer, more responsible LLM usage. Here, we introduced CoCoA, a new framework that enhances uncertainty quantification in LLMs. While CoCoA improves reliability, its use requires careful domain adaptation to align with specific task requirements.
Appendix A Decoding Strategy and Sample Selection
Modern LLMs are capable of producing output using a wide range of decoding strategies, and it is not readily apparent which one to use as a foundation for UQ experiments. On top of that, when sampling multiple outputs stochastically, one has to decide which sample to select for comparison with the target sequence and UQ purposes.
To facilitate the choice of decoding and sample selection strategies for our experiments, we conducted an evaluation of model performance with different approaches to both. Table 3 shows average values of corresponding quality metrics for all combinations of models and datasets. We considered 4 approaches for the selection of output that subsequently is used to calculate the quality of generation.
  • Greedy decoding produces single output by selecting top-1 candidate token at each generation step, thus not further selection of sample is needed.
  • Random sample corresponds to the case where random output is selected among the number of samples produced by repeatedly prompting the model with the same question. In practice we use first generated sample, highlighting model performance when stochastic decoding is done only once.
  • Best (normalized) sample selects the output with highest model-assigned (length-normalized) probability among several sampled outputs.
We note that selecting a random sample from the model outputs incurs a significant drop in the quality of results on several datasets, most prominently on GSM8k. Based on these observations, we evaluate the efficacy of UE on two setups: greedy decoding and stochastic sampling with a focus on the highest-probability sample.
In all experiments, we performed stochastic sampling with temperature $t=1.0$, top-k equal to 50, and top-p equal to $1.0$.
Dataset Metric Llama Mistral Falcon
Greedy Random Best Best Norm Greedy Random Best Best Norm Greedy Random Best Best Norm
Trivia Algin Score 0.686 0.625 0.687 0.689 0.743 0.655 0.750 0.751 0.557 0.473 0.568 0.562
MMLU Accuracy 0.590 0.430 0.597 0.587 0.633 0.558 0.632 0.632 0.713 0.639 0.713 0.713
CoQa Algin Score 0.499 0.359 0.529 0.462 0.574 0.403 0.591 0.528 0.512 0.408 0.560 0.471
GSM8k Accuracy 0.548 0.234 0.261 0.432 0.382 0.169 0.190 0.290 0.776 0.313 0.205 0.593
Xsum Align Score 0.848 0.608 0.825 0.830 0.803 0.578 0.775 0.777 0.842 0.734 0.782 0.809
WMT14FrEn Comet 0.863 0.819 0.852 0.859 0.863 0.812 0.830 0.855 0.867 0.833 0.857 0.863
WMT19DeEn Comet 0.870 0.816 0.854 0.860 0.864 0.805 0.836 0.851 0.846 0.807 0.826 0.838
Table 3: Base quality metrics for models for different evaluated sequence choice.
Appendix B Ablation
B.1 Average Dissimilarity as Uncertainty Measure
Tables 4 and 5 report PRRs of CoCoA-family methods with uncertainty estimates based solely on average dissimilarity of samples, as proposed in equation (8) We observe that it is still widely beneficial to synthesize consistency of outputs with model confidence, even when limiting consistency evaluation to the particular sample to be scored.
Method Dataset
XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
Mistral7b-Base
AveDissimilarity 0.051 0.285 0.500 0.379 0.647 0.423 0.435
$\text{CoCoA}_{MSP}$ 0.330 0.396 0.598 0.383 0.670 0.466 0.517
$\text{CoCoA}_{PPL}$ 0.286 0.375 0.568 0.369 0.674 0.466 0.467
$\text{CoCoA}_{MTE}$ 0.288 0.374 0.564 0.355 0.673 0.447 0.491
Llama8b-Base
AveDissimilarity 0.024 0.389 0.453 0.375 0.614 0.392 0.368
$\text{CoCoA}_{MSP}$ 0.378 0.456 0.582 0.349 0.597 0.485 0.372
$\text{CoCoA}_{PPL}$ 0.387 0.448 0.514 0.338 0.593 0.452 0.433
$\text{CoCoA}_{MTE}$ 0.380 0.446 0.511 0.337 0.601 0.402 0.447
Falcon7b-Base
AveDissimilarity 0.226 0.337 0.496 0.408 0.656 0.485 0.426
$\text{CoCoA}_{MSP}$ 0.257 0.433 0.578 0.396 0.684 0.529 0.436
$\text{CoCoA}_{PPL}$ 0.229 0.436 0.580 0.406 0.677 0.529 0.478
$\text{CoCoA}_{MTE}$ 0.228 0.439 0.577 0.395 0.685 0.517 0.510
Table 4: Comparison of PRRs of CoCoA-family methods with similarity of greedy output and other samples taken in isolation.
Method Dataset
XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
Mistral7b-Base
AveDissimilarity 0.071 0.670 0.708 0.405 0.614 0.423 0.846
$\text{CoCoA}_{MSP}$ 0.366 0.712 0.730 0.430 0.644 0.466 0.562
$\text{CoCoA}_{PPL}$ 0.372 0.735 0.755 0.402 0.648 0.466 0.937
$\text{CoCoA}_{MTE}$ 0.373 0.708 0.732 0.373 0.645 0.447 0.935
Llama8b-Base
AveDissimilarity 0.030 0.473 0.598 0.395 0.600 0.353 0.795
$\text{CoCoA}_{MSP}$ 0.484 0.529 0.685 0.384 0.587 0.452 0.513
$\text{CoCoA}_{PPL}$ 0.450 0.544 0.689 0.364 0.573 0.422 0.925
$\text{CoCoA}_{MTE}$ 0.451 0.520 0.638 0.346 0.582 0.363 0.900
Falcon7b-Base
AveDissimilarity 0.282 0.491 0.651 0.416 0.627 0.484 0.979
$\text{CoCoA}_{MSP}$ 0.210 0.564 0.698 0.428 0.659 0.530 0.498
$\text{CoCoA}_{PPL}$ 0.310 0.579 0.717 0.415 0.644 0.530 1.000
$\text{CoCoA}_{MTE}$ 0.289 0.551 0.678 0.402 0.646 0.517 0.998
Table 5: Comparison of PRRs of CoCoA-family methods with similarity of samples with best sample taken in isolation.
B.2 Choice of Similarity Function
For sample consistency estimation, one could come up with a variety of similarity functions $g(\mathbf{y},\mathbf{y}^{\prime})$. We perform a comparison of the effectiveness of CoCoA-family methods using several such functions. We consider the following functions:
  • AlignScore (Zha et al., 2023) with AlignScore-large model
  • RougeL (Lin, 2004)
  • NLI (He et al., 2021) based on microsoft/deberta-large-mnli model
  • CrossEncoder (Liu et al., 2019) based on cross-encoder/stsb-roberta-large model.
Tables 6 and 7 report these results. There exists a considerable variation of relative effectiveness of proposed methods with various similarity function choices, depending on a task at hand. We opt to report all results in other sections with CrossEncoder-based similarity as it by itself provides a significant improvement over baselines, and for consistency and ease of comparison reasons. However, when applying these methods to a particular task, we encourage users to select appropriate underlying similarity function for best results.
Method Dataset
XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
Mistral7b-Base
$\text{CoCoA}_{MSP}$
AlignScore 0.334 0.293 0.445 0.354 0.655 0.466 0.550
RougeL 0.289 0.358 0.546 0.369 0.649 0.466 0.536
NLI 0.308 0.313 0.477 0.403 0.677 0.470 0.568
CrossEncoder 0.327 0.397 0.595 0.381 0.671 0.466 0.505
$\text{CoCoA}_{PPL}$
AlignScore 0.307 0.308 0.489 0.373 0.666 0.466 0.536
RougeL 0.226 0.369 0.531 0.352 0.653 0.466 0.466
NLI 0.233 0.316 0.501 0.376 0.682 0.470 0.480
CrossEncoder 0.281 0.371 0.565 0.365 0.674 0.466 0.465
$\text{CoCoA}_{MTE}$
AlignScore 0.302 0.299 0.477 0.366 0.664 0.450 0.555
RougeL 0.212 0.377 0.528 0.345 0.652 0.449 0.497
NLI 0.219 0.313 0.488 0.362 0.681 0.453 0.490
CrossEncoder 0.282 0.368 0.560 0.351 0.673 0.448 0.486
Llama8b-Base
$\text{CoCoA}_{MSP}$
AlignScore 0.367 0.331 0.452 0.308 0.596 0.484 0.401
RougeL 0.336 0.393 0.545 0.321 0.563 0.474 0.375
NLI 0.344 0.352 0.467 0.364 0.606 0.478 0.419
CrossEncoder 0.375 0.454 0.583 0.350 0.598 0.480 0.367
$\text{CoCoA}_{PPL}$
AlignScore 0.422 0.346 0.450 0.337 0.596 0.453 0.446
RougeL 0.370 0.408 0.486 0.319 0.552 0.441 0.418
NLI 0.374 0.354 0.438 0.348 0.600 0.446 0.409
CrossEncoder 0.380 0.444 0.514 0.339 0.593 0.447 0.429
$\text{CoCoA}_{MTE}$
AlignScore 0.419 0.340 0.438 0.339 0.605 0.411 0.459
RougeL 0.362 0.417 0.481 0.319 0.560 0.390 0.440
NLI 0.366 0.342 0.428 0.340 0.612 0.396 0.420
CrossEncoder 0.374 0.441 0.511 0.337 0.601 0.394 0.444
Falcon7b-Base
$\text{CoCoA}_{MSP}$
AlignScore 0.278 0.306 0.475 0.361 0.677 0.528 0.470
RougeL 0.205 0.394 0.499 0.378 0.678 0.527 0.417
NLI 0.236 0.361 0.511 0.407 0.684 0.532 0.532
CrossEncoder 0.253 0.436 0.577 0.396 0.685 0.529 0.428
$\text{CoCoA}_{PPL}$
AlignScore 0.252 0.340 0.523 0.410 0.678 0.528 0.521
RougeL 0.170 0.409 0.537 0.389 0.668 0.527 0.439
NLI 0.193 0.364 0.531 0.408 0.680 0.532 0.499
CrossEncoder 0.226 0.437 0.579 0.405 0.677 0.529 0.474
$\text{CoCoA}_{MTE}$
AlignScore 0.253 0.337 0.519 0.403 0.683 0.515 0.554
RougeL 0.170 0.426 0.540 0.382 0.673 0.514 0.472
NLI 0.190 0.364 0.525 0.398 0.687 0.521 0.514
CrossEncoder 0.223 0.438 0.575 0.395 0.685 0.517 0.505
Table 6: Comparison of PRRs of CoCoA-family methods with different choices of similarity function with greedy sample taken in isolation.
Method Dataset
XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
Mistral7b-Base
$\text{CoCoA}_{MSP}$
AlignScore 0.393 0.448 0.491 0.399 0.626 0.467 0.476
RougeL 0.344 0.602 0.597 0.420 0.622 0.466 0.538
NLI 0.340 0.615 0.604 0.445 0.651 0.470 0.456
CrossEncoder 0.366 0.712 0.730 0.430 0.644 0.466 0.562
$\text{CoCoA}_{PPL}$
AlignScore 0.474 0.619 0.657 0.408 0.638 0.467 0.910
RougeL 0.362 0.710 0.717 0.391 0.627 0.466 0.950
NLI 0.370 0.677 0.684 0.414 0.657 0.470 0.941
CrossEncoder 0.372 0.735 0.755 0.402 0.648 0.466 0.937
$\text{CoCoA}_{MTE}$
AlignScore 0.492 0.547 0.590 0.383 0.633 0.449 0.914
RougeL 0.355 0.695 0.684 0.366 0.624 0.448 0.959
NLI 0.364 0.656 0.658 0.387 0.656 0.453 0.918
CrossEncoder 0.373 0.708 0.732 0.373 0.645 0.447 0.935
Llama8b-Base
$\text{CoCoA}_{MSP}$
AlignScore 0.520 0.332 0.491 0.354 0.587 0.457 0.401
RougeL 0.471 0.470 0.588 0.362 0.551 0.446 0.499
NLI 0.466 0.442 0.577 0.386 0.597 0.446 0.470
CrossEncoder 0.484 0.529 0.685 0.384 0.587 0.452 0.513
$\text{CoCoA}_{PPL}$
AlignScore 0.546 0.406 0.561 0.376 0.577 0.429 0.875
RougeL 0.452 0.518 0.639 0.352 0.532 0.417 0.931
NLI 0.458 0.466 0.597 0.365 0.583 0.418 0.912
CrossEncoder 0.450 0.544 0.689 0.364 0.573 0.422 0.925
$\text{CoCoA}_{MTE}$
AlignScore 0.561 0.325 0.497 0.365 0.589 0.380 0.821
RougeL 0.448 0.496 0.598 0.336 0.539 0.361 0.921
NLI 0.449 0.446 0.565 0.344 0.598 0.359 0.881
CrossEncoder 0.451 0.520 0.638 0.346 0.582 0.363 0.900
Falcon7b-Base
$\text{CoCoA}_{MSP}$
AlignScore 0.181 0.378 0.473 0.410 0.654 0.528 0.239
RougeL 0.122 0.531 0.581 0.420 0.655 0.528 0.426
NLI 0.120 0.496 0.607 0.437 0.658 0.533 0.458
CrossEncoder 0.210 0.564 0.698 0.428 0.659 0.530 0.498
$\text{CoCoA}_{PPL}$
AlignScore 0.384 0.454 0.586 0.440 0.648 0.528 0.994
RougeL 0.280 0.565 0.668 0.410 0.637 0.528 1.000
NLI 0.283 0.515 0.671 0.424 0.647 0.533 0.998
CrossEncoder 0.310 0.579 0.717 0.415 0.644 0.530 1.000
$\text{CoCoA}_{MTE}$
AlignScore 0.292 0.386 0.498 0.435 0.648 0.515 0.972
RougeL 0.222 0.545 0.607 0.400 0.633 0.515 0.998
NLI 0.201 0.498 0.636 0.415 0.645 0.521 0.987
CrossEncoder 0.289 0.551 0.678 0.402 0.646 0.517 0.998
Table 7: Comparison of PRRs of CoCoA-family methods with different choices of similarity function with best sample taken in isolation.
B.3 Different Ways of Combining Confidence and Consistency
We justify the particular form of equation (9) by considering alternative ways to combine sample-focused confidence with consistency estimation. Results are presented in Tables 8 and 9. In particular, we investigate the performance of the additive approach (AdditiveCoCoA):
$\mathrm{U}_{AdditiveCoCoA}=u_{*}^{\text{info}}+u_{*}^{\text{cons}},$ (11)
and the same multiplicative combination, replacing sample-focused dissimilarity from (8) with the average of the full pairwise dissimilarity matrix (19):
$\mathrm{U}_{FullSampleCoCoA}=u_{*}^{\text{info}}\cdot U_{Deg}.$ (12)
It is evident that on average the multiplicative form proposed in equation (9) with both confidence and consistency terms focused on a single sample is the better performing variant.
Method Dataset
XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
Mistral7b-Base
$\text{AdditiveCoCoA}_{MSP}$ 0.290 0.319 0.459 0.351 0.654 0.471 0.472
$\text{FullSampleCoCoA}_{MSP}$ 0.319 0.385 0.590 0.357 0.668 0.467 0.505
$\text{ProbCoCoA}_{MSP}$ 0.059 0.302 0.520 0.390 0.671 0.461 0.435
$\text{CoCoA}_{MSP}$ 0.330 0.396 0.598 0.383 0.670 0.466 0.517
$\text{AdditiveCoCoA}_{PPL}$ 0.262 0.392 0.564 0.369 0.671 0.464 0.494
$\text{FullSampleCoCoA}_{PPL}$ 0.277 0.373 0.551 0.334 0.672 0.467 0.435
$\text{ProbCoCoA}_{PPL}$ 0.297 0.369 0.566 0.373 0.674 0.464 0.475
$\text{CoCoA}_{PPL}$ 0.286 0.375 0.568 0.369 0.674 0.466 0.467
$\text{AdditiveCoCoA}_{MTE}$0.279 -0.058 -0.072 0.098 0.312 0.079 0.187
$\text{FullSampleCoCoA}_{MTE}$ 0.274 0.368 0.543 0.309 0.668 0.442 0.456
$\text{CoCoA}_{MTE}$ 0.288 0.374 0.564 0.355 0.673 0.447 0.491
Llama8b-Base
$\text{AdditiveCoCoA}_{MSP}$ 0.330 0.345 0.462 0.301 0.566 0.502 0.326
$\text{FullSampleCoCoA}_{MSP}$ 0.358 0.434 0.564 0.333 0.589 0.488 0.354
$\text{ProbCoCoA}_{MSP}$ 0.031 0.405 0.471 0.371 0.612 0.461 0.368
$\text{CoCoA}_{MSP}$ 0.378 0.456 0.582 0.349 0.597 0.485 0.372
$\text{AdditiveCoCoA}_{PPL}$ 0.368 0.431 0.504 0.336 0.595 0.455 0.437
$\text{FullSampleCoCoA}_{PPL}$ 0.389 0.420 0.487 0.314 0.580 0.450 0.399
$\text{ProbCoCoA}_{PPL}$ 0.381 0.445 0.513 0.345 0.599 0.446 0.438
$\text{CoCoA}_{PPL}$ 0.387 0.448 0.514 0.338 0.593 0.452 0.433
$\text{AdditiveCoCoA}_{MTE}$0.331 -0.042 -0.122 0.089 0.321 -0.122 0.117
$\text{FullSampleCoCoA}_{MTE}$ 0.383 0.410 0.481 0.308 0.588 0.363 0.414
$\text{CoCoA}_{MTE}$ 0.380 0.446 0.511 0.337 0.601 0.402 0.447
Falcon7b-Base
$\text{AdditiveCoCoA}_{MSP}$ 0.203 0.318 0.409 0.350 0.674 0.533 0.379
$\text{FullSampleCoCoA}_{MSP}$ 0.225 0.423 0.571 0.388 0.678 0.533 0.404
$\text{ProbCoCoA}_{MSP}$ 0.226 0.367 0.515 0.416 0.680 0.526 0.426
$\text{CoCoA}_{MSP}$ 0.257 0.433 0.578 0.396 0.684 0.529 0.436
$\text{AdditiveCoCoA}_{PPL}$ 0.222 0.433 0.580 0.413 0.677 0.525 0.489
$\text{FullSampleCoCoA}_{PPL}$ 0.204 0.425 0.565 0.393 0.669 0.533 0.437
$\text{ProbCoCoA}_{PPL}$ 0.235 0.433 0.576 0.410 0.680 0.528 0.482
$\text{CoCoA}_{PPL}$ 0.229 0.436 0.580 0.406 0.677 0.529 0.478
$\text{AdditiveCoCoA}_{MTE}$ 0.001 -0.103 -0.106 0.114 0.041 0.138 0.221
$\text{FullSampleCoCoA}_{MTE}$ 0.201 0.425 0.557 0.377 0.675 0.519 0.470
$\text{CoCoA}_{MTE}$ 0.228 0.439 0.577 0.395 0.685 0.517 0.510
Table 8: Comparison of PRRs of CoCoA-family methods with alternative formulations with greedy sample taken in isolation.
Method Dataset
XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
Mistral7b-Base
$\text{AdditiveCoCoA}_{MSP}$ 0.333 0.239 0.310 0.406 0.631 0.472 0.311
$\text{FullSampleCoCoA}_{MSP}$ 0.354 0.543 0.565 0.412 0.643 0.468 0.428
$\text{ProbCoCoA}_{MSP}$ 0.076 0.684 0.721 0.428 0.643 0.464 0.846
$\text{CoCoA}_{MSP}$ 0.366 0.712 0.730 0.430 0.644 0.466 0.562
$\text{AdditiveCoCoA}_{PPL}$ 0.368 0.737 0.751 0.406 0.644 0.465 0.939
$\text{FullSampleCoCoA}_{PPL}$ 0.383 0.714 0.723 0.379 0.649 0.468 0.933
$\text{ProbCoCoA}_{PPL}$ 0.369 0.738 0.756 0.401 0.649 0.467 0.935
$\text{CoCoA}_{PPL}$ 0.372 0.735 0.755 0.402 0.648 0.466 0.937
$\text{AdditiveCoCoA}_{MTE}$ 0.368 0.723 0.702 0.332 0.643 0.452 0.942
$\text{FullSampleCoCoA}_{MTE}$ 0.380 0.661 0.653 0.331 0.643 0.442 0.929
$\text{CoCoA}_{MTE}$ 0.373 0.708 0.732 0.373 0.645 0.447 0.935
Llama8b-Base
$\text{AdditiveCoCoA}_{MSP}$ 0.466 0.349 0.425 0.333 0.555 0.473 0.285
$\text{FullSampleCoCoA}_{MSP}$ 0.476 0.462 0.619 0.363 0.574 0.464 0.379
$\text{ProbCoCoA}_{MSP}$ 0.035 0.491 0.617 0.398 0.598 0.433 0.795
$\text{CoCoA}_{MSP}$ 0.484 0.529 0.685 0.384 0.587 0.452 0.513
$\text{AdditiveCoCoA}_{PPL}$ 0.454 0.536 0.673 0.358 0.575 0.425 0.923
$\text{FullSampleCoCoA}_{PPL}$ 0.459 0.525 0.649 0.343 0.556 0.430 0.914
$\text{ProbCoCoA}_{PPL}$ 0.438 0.547 0.689 0.364 0.574 0.419 0.923
$\text{CoCoA}_{PPL}$ 0.450 0.544 0.689 0.364 0.573 0.422 0.925
$\text{AdditiveCoCoA}_{MTE}$ 0.457 0.496 0.579 0.304 0.561 0.361 0.901
$\text{FullSampleCoCoA}_{MTE}$ 0.455 0.464 0.577 0.313 0.563 0.341 0.878
$\text{CoCoA}_{MTE}$ 0.451 0.520 0.638 0.346 0.582 0.363 0.900
Falcon7b-Base
$\text{AdditiveCoCoA}_{MSP}$ 0.100 0.397 0.394 0.393 0.649 0.534 -0.156
$\text{FullSampleCoCoA}_{MSP}$ 0.144 0.531 0.607 0.416 0.654 0.533 0.189
$\text{ProbCoCoA}_{MSP}$ 0.282 0.522 0.670 0.434 0.658 0.529 0.978
$\text{CoCoA}_{MSP}$ 0.210 0.564 0.698 0.428 0.659 0.530 0.498
$\text{AdditiveCoCoA}_{PPL}$ 0.297 0.582 0.706 0.417 0.643 0.526 1.000
$\text{FullSampleCoCoA}_{PPL}$ 0.297 0.560 0.670 0.405 0.641 0.533 1.000
$\text{ProbCoCoA}_{PPL}$ 0.311 0.587 0.718 0.414 0.648 0.531 1.000
$\text{CoCoA}_{PPL}$ 0.310 0.579 0.717 0.415 0.644 0.530 1.000
$\text{AdditiveCoCoA}_{MTE}$ 0.253 0.554 0.634 0.383 0.630 0.523 0.997
$\text{FullSampleCoCoA}_{MTE}$ 0.237 0.502 0.554 0.383 0.636 0.519 0.989
$\text{CoCoA}_{MTE}$ 0.289 0.551 0.678 0.402 0.646 0.517 0.998
Table 9: Comparison of PRRs of CoCoA-family methods with alternative formulations of with best sample taken in isolation.
Appendix C Detailed Description of Uncertainty Quantification Methods
In this section, we provide a detailed description of the uncertainty quantification methods used in this study.
C.1 Information-Based Methods
Information-based methods are commonly used to estimate uncertainty by analyzing the probability distributions of tokens within a given output. These methods examine different levels of model generation, such as the model’s confidence in producing a specific sequence, its ability to predict individual tokens at each generation step, and the variability in the token-level predictions across the sequence.
Maximum Sequence Probability (MSP) is one of the simplest and most direct methods for estimating uncertainty. It measures the probability of the most likely output sequence given a specific input. Thus, uncertainty is quantified by calculating the probability of the sequence with the highest likelihood, under the assumption that the model is most confident in this output. It is defined as:
$U_{MSP}(\mathbf{y}\mid\mathbf{x},\bm{\theta})=-\log P(\mathbf{y}\mid\mathbf{x}).$ (13)
Perplexity (PPL) is another widely used metric for estimating uncertainty in language models (Fomicheva et al., 2020). It measures the model’s confidence by evaluating the average likelihood of generating the sequence tokens:
$U_{\mathrm{PPL}}(\mathbf{y},\mathbf{x})=-\frac{1}{L}\log P(\mathbf{y}\mid\mathbf{x}).$ (14)
Mean Token Entropy takes a broader view of uncertainty by considering the token-level predictions across the entire sequence (Fomicheva et al., 2020). Instead of evaluating the model’s confidence in a single output or individual token predictions, Mean Token Entropy calculates the average entropy of the token probability distributions for each token in the sequence:
$U_{\mathcal{H}_{T}}(\mathbf{y},\mathbf{x})=\frac{1}{L}\sum_{l=1}^{L}\mathcal{H}(y_{l}\mid\mathbf{y}_{<l},\mathbf{x}),$ (15)
where $\mathcal{H}(y_{l}\mid\mathbf{y}_{<l},\mathbf{x})$ is an entropy of the token distribution $P(y_{l}\mid\mathbf{y}_{<l},\mathbf{x})$.
The TokenSAR method, introduced in (Duan et al., 2024), generalizes length-normalized log probability by computing a weighted average of the negative log probabilities of generated tokens, where weights are based on token relevance to the overall text. Using a similarity function $g(\cdot,\cdot)$ and token relevance function $R_{T}(y_{k},\mathbf{y},\mathbf{x})=1-g(\mathbf{x}\cup\mathbf{y},\mathbf{x}\cup\mathbf{y}\setminus y_{k})$, the uncertainty estimate is calculated as:
$U_{\mathrm{TokenSAR}}(\mathbf{y},\mathbf{x})=-\sum_{l=1}^{L}\tilde{\mathrm{R}}_{T}(y_{l},\mathbf{y},\mathbf{x})\log P(y_{l}\mid\mathbf{y}_{<l},\mathbf{x}),$ (16)
where
$\tilde{\mathrm{R}}_{T}(y_{k},\mathbf{y},\mathbf{x})=\frac{\mathrm{R}_{T}(y_{k},\mathbf{y},\mathbf{x})}{\sum\nolimits_{l=1}^{L}\mathrm{R}_{T}(y_{l},\mathbf{y},\mathbf{x})}.$ (17)
This measure is central for computing SAR uncertainty measure.
C.2 Consistency-Based Methods
Consistency-based methods assess the uncertainty of a language model by evaluating the semantic consistency of its predictions across multiple outputs for the same prompt. The core idea is that semantically similar outputs indicate higher confidence, while diverse or conflicting outputs suggest greater uncertainty. Since language models can express the same meaning in different surface forms, these methods construct a semantic similarity matrix $G=(g_{ij})$, where each entry represents the degree of similarity between pairs of responses. By clustering responses into groups with equivalent meanings, these methods provide a semantic measure of the model’s consistency.
Lin et al. (2024) offers two similarity measures to evaluate the similarity of sequences. The first is the Jaccard similarity, which treats sequences as sets of words and calculates the proportion of shared words to the total number of unique words in both sequences: $g(\mathbf{y},\mathbf{y}^{\prime})=|\mathbf{y}\cap\mathbf{y}^{\prime}|/|\mathbf{y}\cup\mathbf{y}^{\prime}|$.
Natural Language Inference (NLI) provides another method for computing similarity between sequences. We use the DeBERTa-large NLI model (He et al., 2021), following Kuhn et al. (2023). For each pair of sequences, an NLI model predicts two probabilities: ${p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})$, indicating entailment, and ${p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$, indicating contradiction. Similarity is then defined as either $g_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})={p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})$ or $g_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})=1-{p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$.
Among the simplest consistency-based approaches are the Number of Semantic Sets and the Sum of Eigenvalues of the Graph Laplacian (Lin et al., 2024). Number of Semantic Sets estimates how many distinct “meanings” the model produces by clustering its outputs with an NLI model. The number of semantic sets is initially equal to the total number of generated answers, $M$. Two sentences are grouped into the same cluster if the following conditions are satisfied: ${p}_{\text{entail}}(\mathbf{y}^{i},\mathbf{y}^{j})>{p}_{\text{contra}}(\mathbf{y}^{i},\mathbf{y}^{j})\quad\text{and}\quad{p}_{\text{entail}}(\mathbf{y}^{j},\mathbf{y}^{i})>{p}_{\text{contra}}(\mathbf{y}^{j},\mathbf{y}^{i})$. This computation is performed for all pairs of answers, and the final number of distinct clusters is denoted by $U_{\text{NumSemSets}}$.
Sum of Eigenvalues of the Graph Laplacian examines global diversity: it constructs a similarity matrix among the sampled outputs and computes a continuous uncertainty score from the eigenvalues of the Laplacian of that similarity graph. Lin et al. (2024) proposes computing an averaged similarity matrix as $g_{ij}=\bigl{(}g\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}+g\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(i)}\bigr{)}\bigr{)}/2$. The Laplacian for the matrix $G$ is defined as $L=I-D^{-\frac{1}{2}}GD^{-\frac{1}{2}}$, where $D$ is a diagonal matrix with elements $D_{ii}=\sum_{j=1}^{M}g_{ij}$. Consequently, the following formula is derived:
$U_{\text{EigV}}=\sum_{i=1}^{M}\max(0,1-\lambda_{i}).$ (18)
Both Number of Semantic Sets and Sum of Eigenvalues of the Graph Laplacian effectively capture overall variation in generated text but cannot produce an individual uncertainty score for each output. To address this, Lin et al. (2024) proposes to use the diagonal Degree Matrix $D$ which represents the total similarity of each answer with all others. The corrected trace of $D$ provides an average pairwise distance between answers, and uncertainty is computed as:
$U_{\mathrm{DegMat}}=1-\mathrm{trace}(D)/M^{2}.$ (19)
C.3 Information-Based Methods with Repeated Sampling
In this section we detail methods that integrate model confidence with consistency.
We can compute the entropy on the sequence level $\mathbb{E}\bigl{[}-\log P(\mathbf{y}\mid\mathbf{x})\bigr{]}$, where the expectation is taken over the sequences $\mathbf{y}$ randomly generated from the distribution $P(\mathbf{y}\mid\mathbf{x})$. Unfortunately, while for token level, we have an exact way of computing the entropy, for the sequence level, we need to adhere to some approximations. In practice, we can use Monte-Carlo integration, i.e. generate several sequences $\mathbf{y}^{(i)},\,i=1,\dots,M$ via random sampling and compute Monte Carlo Sequence Entropy:
$U_{\mathcal{H}_{S}}(\mathbf{x})=-\frac{1}{M}\sum_{i=1}^{M}\log P(\mathbf{y}^{(i)}\mid\mathbf{x}).$ (20)
We can replace $P(\mathbf{y}^{(i)}\mid\mathbf{x})$ with its length-normalized version $\bar{P}(\mathbf{y}^{(i)}\mid\mathbf{x})$ leading to a more reliable uncertainty measure in some cases.
Semantic Entropy (Kuhn et al., 2023) addresses the issue of generated sequences with similar meanings but differing probabilities according to the model, which can heavily influence the resulting entropy value (20). The method clusters generated sequences $\mathbf{y}^{(i)},\,i=1,\dots,M$ into semantically homogeneous groups $\mathcal{C}_{k},~{}k=1,\dots,K$ (where $K\leq M$) using a bi-directional entailment algorithm. Probabilities of sequences are averaged within each cluster. The entropy estimate is then defined as:
$U_{\mathrm{SE}}(\mathbf{x})=-\sum_{k=1}^{K}\frac{|\mathcal{C}_{k}|}{M}\log\hat{P}_{k}(\mathbf{x}),$ (21)
where $\hat{P}_{k}(\mathbf{x})=\sum_{\mathbf{y}\in\mathcal{C}_{k}}P(\mathbf{y}\mid\mathbf{x})$ represents the aggregated probability for cluster $\mathcal{C}_{k}$.
SentenceSAR (Duan et al., 2024) enhances the probability of sentences that are more relevant. It uses a sentence relevance measure $g\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(k)}\bigr{)}$ to evaluate the relevance of $\mathbf{y}^{(j)}$ with respect to $\mathbf{y}^{(k)}$. SentenceSAR is calculated as:
$U_{\mathrm{SentSAR}}(\mathbf{x})=-\frac{1}{M}\sum_{i=1}^{M}\log\Bigl{(}P(\mathbf{y}^{(i)}\mid\mathbf{x})+\frac{1}{t}\mathrm{R}_{S}(\mathbf{y}^{(i)},\mathbf{x})\Bigr{)},$ (22)
where $t$ is a temperature parameter used to control the scale of shifting to relevance, and
$\mathrm{R}_{S}(\mathbf{y}^{(j)},\mathbf{x})\!=\sum_{k\neq j}g\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(k)}\bigr{)}P\bigl{(}\mathbf{y}^{(k)}\mid\mathbf{x}\bigr{)}.$ (23)
The combination of SentenceSAR and TokenSAR results in a unified method called SAR (Duan et al., 2024). In this approach, the generative probability $P(\mathbf{y}\mid\mathbf{x})$ in the SentenceSAR formula is replaced with the token-shifted probability $P^{\prime}(\mathbf{y}\mid\mathbf{x})=\exp\bigl{\{}-\mathrm{TokenSAR}(\mathbf{y},\mathbf{x})\bigr{\}}$, creating a comprehensive measure that integrates both sentence- and token-level adjustments.
Appendix D Detailed Experimental Results
In this section, we present detailed experimental results, which were used for computing values in Tables 1 and 2.
Method Dataset
XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
Mistral7b-Base
MCSE 0.007 0.257 0.350 0.247 0.496 0.337 0.475
MCNSE 0.009 0.342 0.478 0.238 0.540 0.356 0.401
Semantic Entropy 0.008 0.271 0.382 0.271 0.562 0.387 0.472
SAR 0.094 0.353 0.517 0.313 0.644 0.419 0.471
DegMat 0.137 0.229 0.382 0.336 0.646 0.410 0.299
EigValLaplacian 0.132 0.207 0.328 0.301 0.624 0.398 0.241
MSP 0.287 0.315 0.451 0.326 0.628 0.474 0.471
$\text{CoCoA}_{MSP}$ 0.330 0.396 0.598 0.383 0.670 0.466 0.517
PPL 0.204 0.365 0.489 0.281 0.632 0.474 0.311
$\text{CoCoA}_{PPL}$ 0.286 0.375 0.568 0.369 0.674 0.466 0.467
MTE 0.182 0.392 0.484 0.243 0.619 0.456 0.350
$\text{CoCoA}_{MTE}$ 0.288 0.374 0.564 0.355 0.673 0.447 0.491
Llama8b-Base
MCSE 0.033 0.293 0.354 0.237 0.482 0.171 0.351
MCNSE 0.022 0.370 0.415 0.219 0.501 0.170 0.344
Semantic Entropy 0.033 0.297 0.389 0.272 0.549 0.229 0.375
SAR 0.077 0.427 0.483 0.311 0.595 0.352 0.398
DegMat 0.081 0.250 0.355 0.353 0.622 0.342 0.309
EigValLaplacian 0.079 0.198 0.278 0.332 0.604 0.292 0.273
MSP 0.328 0.342 0.456 0.277 0.526 0.508 0.324
$\text{CoCoA}_{MSP}$ 0.378 0.456 0.582 0.349 0.597 0.485 0.372
PPL 0.369 0.351 0.422 0.253 0.507 0.461 0.303
$\text{CoCoA}_{PPL}$ 0.387 0.448 0.514 0.338 0.593 0.452 0.433
MTE 0.357 0.357 0.408 0.239 0.497 0.350 0.326
$\text{CoCoA}_{MTE}$ 0.380 0.446 0.511 0.337 0.601 0.402 0.447
Falcon7b-Base
MCSE 0.159 0.297 0.337 0.258 0.549 0.420 0.427
MCNSE 0.108 0.371 0.474 0.293 0.586 0.442 0.299
Semantic Entropy 0.164 0.307 0.389 0.294 0.581 0.463 0.418
SAR 0.144 0.398 0.517 0.381 0.649 0.508 0.387
DegMat 0.201 0.274 0.431 0.407 0.651 0.480 0.395
EigValLaplacian 0.201 0.229 0.394 0.381 0.645 0.454 0.358
MSP 0.201 0.312 0.400 0.321 0.662 0.539 0.377
$\text{CoCoA}_{MSP}$ 0.257 0.433 0.578 0.396 0.684 0.529 0.436
PPL 0.155 0.375 0.525 0.316 0.644 0.539 0.326
$\text{CoCoA}_{PPL}$ 0.229 0.436 0.580 0.406 0.677 0.529 0.478
MTE 0.152 0.409 0.537 0.291 0.633 0.533 0.367
$\text{CoCoA}_{MTE}$ 0.228 0.439 0.577 0.395 0.685 0.517 0.510
Table 10: Detailed experimental results with greedy sample taken in isolation.
Method Dataset
XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
Mistral7b-Base
MCSE 0.162 0.406 0.407 0.289 0.492 0.339 0.693
MCNSE 0.196 0.471 0.507 0.277 0.529 0.358 0.700
Semantic Entropy 0.164 0.434 0.442 0.312 0.554 0.389 0.675
SAR 0.175 0.563 0.590 0.347 0.620 0.421 0.780
DegMat 0.205 0.439 0.410 0.376 0.618 0.410 0.454
EigValLaplacian 0.197 0.388 0.344 0.342 0.600 0.399 0.361
MSP 0.330 0.212 0.291 0.388 0.607 0.476 0.307
$\text{CoCoA}_{MSP}$ 0.366 0.712 0.730 0.430 0.644 0.466 0.562
PPL 0.365 0.695 0.676 0.327 0.615 0.476 0.931
$\text{CoCoA}_{PPL}$ 0.372 0.735 0.755 0.402 0.648 0.466 0.937
MTE 0.350 0.668 0.606 0.254 0.594 0.457 0.932
$\text{CoCoA}_{MTE}$ 0.373 0.708 0.732 0.373 0.645 0.447 0.935
Llama8b-Base
MCSE 0.192 0.366 0.395 0.259 0.465 0.158 0.546
MCNSE 0.186 0.377 0.480 0.239 0.484 0.165 0.634
Semantic Entropy 0.194 0.371 0.451 0.286 0.528 0.213 0.559
SAR 0.159 0.441 0.571 0.327 0.578 0.340 0.668
DegMat 0.191 0.274 0.409 0.367 0.606 0.320 0.397
EigValLaplacian 0.190 0.216 0.333 0.340 0.587 0.274 0.351
MSP 0.464 0.339 0.413 0.304 0.514 0.483 0.281
$\text{CoCoA}_{MSP}$ 0.484 0.529 0.685 0.384 0.587 0.452 0.513
PPL 0.458 0.504 0.622 0.294 0.483 0.441 0.912
$\text{CoCoA}_{PPL}$ 0.450 0.544 0.689 0.364 0.573 0.422 0.925
MTE 0.449 0.437 0.501 0.239 0.458 0.326 0.884
$\text{CoCoA}_{MTE}$ 0.451 0.520 0.638 0.346 0.582 0.363 0.900
Falcon7b-Base
MCSE 0.128 0.399 0.419 0.285 0.535 0.421 0.598
MCNSE 0.153 0.395 0.452 0.318 0.588 0.443 0.771
Semantic Entropy 0.134 0.420 0.460 0.319 0.566 0.463 0.567
SAR 0.193 0.455 0.521 0.385 0.642 0.509 0.826
DegMat 0.177 0.350 0.422 0.422 0.637 0.480 0.633
EigValLaplacian 0.174 0.289 0.382 0.393 0.622 0.454 0.522
MSP 0.099 0.385 0.378 0.369 0.638 0.540 -0.175
$\text{CoCoA}_{MSP}$ 0.210 0.564 0.698 0.428 0.659 0.530 0.498
PPL 0.275 0.541 0.637 0.353 0.614 0.540 1.000
$\text{CoCoA}_{PPL}$ 0.310 0.579 0.717 0.415 0.644 0.530 1.000
MTE 0.186 0.475 0.510 0.317 0.573 0.534 0.984
$\text{CoCoA}_{MTE}$ 0.289 0.551 0.678 0.402 0.646 0.517 0.998
Table 11: Detailed experimental results with best sample taken in isolation.

### v3

@@
-CoCoA: A Generalized Approach to Uncertainty Quantification by Integrating Confidence and Consistency of LLM Outputs
+Uncertainty Quantification for LLMs through Minimum Bayes Risk: Bridging Confidence and Consistency
 Abstract
-Uncertainty quantification (UQ) methods for Large Language Models (LLMs) encompasses a variety of approaches, with two major types being particularly prominent: information-based, which focus on model confidence expressed as token probabilities, and consistency-based, which assess the semantic relationship between multiple outputs generated using repeated sampling.
-Several recent methods have combined these two approaches and shown impressive performance in various applications.
+Uncertainty quantification (UQ) methods for Large Language Models (LLMs) encompass a variety of approaches, with two major types being particularly prominent: information-based, which focus on model confidence expressed as token probabilities, and consistency-based, which assess the semantic relationship between multiple outputs generated using repeated sampling.
+Several recent methods have combined these two approaches to boost UQ performance.
 However, they sometimes fail to outperform much simpler baseline methods.
-Our investigation reveals distinctive characteristics of LLMs as probabilistic models, which help to explain why these UQ methods underperform in certain tasks.
-Based on these findings, we propose a new way of synthesizing model confidence and output consistency that leads to a family of efficient and robust UQ methods.
-We evaluate our approach across a variety of tasks such as question answering, abstractive summarization, and machine translation, demonstrating sizable improvements over state-of-the-art UQ approaches.
+Our work discusses the fundamental approach to constructing uncertainty measures that directly links uncertainty with the minimum Bayes risks achieved by LLM decoding.
+Building on these findings, we propose a novel approach to integrating model confidence with output consistency, resulting in a family of efficient and robust UQ methods.
+We evaluate our approach across various tasks such as question answering, abstractive summarization, and machine translation, demonstrating sizable improvements over state-of-the-art UQ approaches.
 1 Introduction
-Large Language Models (LLMs) have revolutionized natural language processing (NLP), enabling advancements in information retrieval, question answering, machine translation, and other language-driven applications.
-As these models become an integral part of everyday life, ensuring the reliability of their outputs is crucial, especially in high-stakes scenarios where errors or uncertainty can have serious consequences.
-One way to address this challenge is through uncertainty quantification (UQ), which measures how confident a model is in its outputs, and makes possible the rejection of generations with a high risk of being incorrect.
-UQ for LLMs is a rapidly advancing research area, with new methods for estimating uncertainty emerging each year.
-The large portion of novel techniques is based on two fundamental approaches: information-theoretic analysis and the assessment of output consistency.
 Figure 1: Example of inconsistent probabilities assigned to semantically identical answers by an LLM, demonstrating the limitation of relying solely on sequence-level information.
+Large Language Models (LLMs) have revolutionized natural language processing (NLP), enabling advances in information retrieval Zhu et al.
+(2023), question answering Kwiatkowski et al.
+(2019), machine translation Kocmi and Federmann (2023), and a broad range of other NLP applications.
+As these models become an integral part of our everyday life, ensuring the reliability of their outputs is crucial, especially in high-stakes scenarios where errors can have serious consequences.
+One way to address this challenge is through uncertainty quantification (UQ), which estimates the reliability of model outputs and enables abstention from generations that are likely to be incorrect.
+UQ for LLMs is a rapidly advancing research area, with new UQ methods emerging each year.
+Most novel techniques are based on two fundamental approaches: (a) information-theoretic analysis or (b) assessment of output consistency.
 Information-theoretic methods quantify the confidence of a model by analyzing the probability distributions it induces for predictions (Malinin and Gales, 2021; Fomicheva et al., 2020).
-A key limitation of these methods is that they cannot account for the semantic variability across multiple possible outputs for the same input.
+A key limitation of these methods is that they cannot account for semantic variability across multiple possible responses for the same input.
 Specifically, the model may generate answers with the same meaning but with very different assigned probabilities; see Figure 1.
 LLMs are trained to predict the next token in a sequence based on patterns observed in vast amounts of data, resulting in varying probabilities for semantically equivalent output sequences.
-In contrast, consistency-based methods directly analyze the semantic relationships between the sampled outputs (Lin et al., 2024; Fomicheva et al., 2020).
-Information-theoretic and consistency-based methods have complementary strengths: the former provides insights into the model’s internal confidence, while the latter captures the uncertainty as objective variability of meaning among sampled outputs.
+In contrast, consistency-based methods directly analyze the semantic relationship between the sampled outputs (Fomicheva et al., 2020; Lin et al., 2024), capturing uncertainty as objective variability of meaning among the sampled outputs.
+Information-theoretic and consistency-based methods have complementary strengths.
 For this reason, recent state-of-the-art methods aimed to unify these approaches (Kuhn et al., 2023; Duan et al., 2024).
-Although such methods show good performance in various applications, they sometimes fail to outperform their simpler counterparts in certain scenarios (Vashurin et al., 2024).
-Our investigation revealed distinctive characteristics of LLMs as probabilistic models, shedding light on why current state-of-the-art UQ methods that attempt to integrate both approaches often underperform in certain tasks.
-Specifically, we highlight the complexity of the token prediction process and the absence of a unified framework that simultaneously addresses model confidence and output variability, both of which can limit the effectiveness of existing UQ techniques.
-This insight drives our proposal for a novel family of methods that integrate model confidence with output consistency, resulting in more efficient and robust UQ techniques.
+We follow this direction and propose a way of quantifying risk as a combination of these basic measures.
+This results in a family of efficient and robust UQ techniques.
 Our approach combines the strengths of both information-based and consistency-based methods, providing a more comprehensive and accurate assessment of uncertainty.
-Our main contributions can be summarized as follows.
-  • We identify key limitations in current UQ methods for LLMs, particularly in addressing both token- and sequence-level confidence and output consistency.
-  • We present a family of Confidence and Consistency-based Approaches (CoCoA) to UQ, offering a new way to merge information- and consistency-based measures for uncertainty quantification in LLMs.
-  • We evaluate our approaches across a variety of NLP tasks, including question answering, summarization, and translation.
+Recently, it has been argued Wang and Holmes (2025); Daheim et al.
+(2025) that UQ in NLP tasks can be viewed through the lens of the Minimum Bayes Risk (MBR) framework.
+Following this, we formulate our methods as particular forms of risk functions under the MBR approach.
+Our main contributions can be summarized as follows:
+  • We propose a new way of quantifying risk that combines information-theoretic and consistency-based measures, and derive a family of Confidence and Consistency-based Approaches (CoCoA) to uncertainty quantification in LLMs.
+  • We show that these measures can be approximated by a learned function trained on unlabeled held-out set with a negligible loss of performance, eliminating the need for costly repeated sampling from the LLM.
+  • We evaluate our approaches across a variety of NLP tasks, including question answering, summarization, and machine translation.
 Our experiments demonstrate sizable improvements in the reliability and robustness of UQ compared to state-of-the-art methods.
-Figure 2: Illustration of the method: the LLM generates a response, evaluates the similarity to alternatives, computes the uncertainty, and combines the uncertainty score with the similarity measures.
-High similarity to alternatives reduces the uncertainty, while low similarity keeps it high.
 2 Background
-In this section, we introduce key concepts related to uncertainty quantification for LLMs, outline existing methods, discuss their limitations, and highlight the motivation for our approach.
-First and foremost, it is important to establish the concept of an uncertainty function.
-Let $\mathbf{y}=f(\mathbf{x})$ denote the output of an LLM given an input sequence $\mathbf{x}$.
-The model defines a probabilistic output distribution $p(\mathbf{y}\mid\mathbf{x})$, from which outputs can be sampled.
-An uncertainty function $U$ is a mapping that quantifies the level of uncertainty $u$ associated with the output of a model $\mathbf{y}$, conditioned on the input sequence $\mathbf{x}$, which we denote as
-$u=U(\mathbf{y}\mid\mathbf{x}).$ (1)
-With this notion in mind, we now discuss existing methods for uncertainty quantification, outline their limitations, and describe how our approach addresses them.
-2.1 Single-Sequence Information-Based Methods
-Information-based methods rely on a single sample from the LLM and estimate the uncertainty of the generated sequence by aggregating the uncertainty scores of individual tokens.
-One of the simplest techniques of this kind is Maximum Sequence Probability (MSP):
-$u_{\text{MSP}}=-\log p\bigl{(}\mathbf{y}\mid\mathbf{x}\bigr{)}.$ (2)
-Several other measures fall into this category, including Perplexity and Mean Token Entropy (Fomicheva et al., 2020); see Appendix C.1 for details.
+2.1 Language Model Decoding
+LLMs define a probabilistic output distribution $p(\mathbf{y}\mid\mathbf{x})$ for the given input sequence $\mathbf{x}$.
+The standard ways to obtain a particular output $\mathbf{y}_{*}$ given $p(\mathbf{y}\mid\mathbf{x})$ include various variants of sampling $\mathbf{y}_{*}\sim p(\mathbf{y}\mid\mathbf{x})$ and greedy decoding, where
+$\mathbf{y}_{*}=\arg\max_{\mathbf{y}}p(\mathbf{y}\mid\mathbf{x}).$ (1)
+An alternative approach to decoding is Minimum Bayes Risk (MBR) decoding Kumar and Byrne (2004):
+$\mathbf{y}_{*}=\arg\min_{\mathbf{y}\in\mathcal{Y}}R(\mathbf{y}\mid\mathbf{x}),$ (2)
+where $\mathcal{Y}$ is the set of candidate sequences and $R(\mathbf{y}\mid\mathbf{x})$ is a risk function:
+$R(\mathbf{y}\mid\mathbf{x})=\mathbb{E}_{\mathbf{y}^{\prime}\sim p(\mathbf{y}\mid\mathbf{x})}\,\,r(\mathbf{y},\mathbf{y}^{\prime})$ (3)
+with $r(\mathbf{y},\mathbf{y}^{\prime})$ being a pairwise loss function.
+The standard choice in the MBR literature is to take $r(\mathbf{y},\mathbf{y}^{\prime})\propto-s(\mathbf{y},\mathbf{y}^{\prime})$, where the so-called utility function $s(\mathbf{y},\mathbf{y}^{\prime})$ represents some notion of similarity between generated sequences $\mathbf{y}$ and $\mathbf{y}^{\prime}$.
+Below we discuss how various decoding strategies and corresponding risk functions lead to well-grounded definitions of uncertainty.
+2.2 From Risk to Uncertainty
+Generally, an uncertainty function $U$ is a mapping that quantifies the level of uncertainty associated with the output of a model $\mathbf{y}$, conditioned on the input sequence $\mathbf{x}$, which we denote as $U(\mathbf{y}\mid\mathbf{x})$.
+Bayesian uncertainty measures are well known to be strongly connected with minimum risks of various kinds Xu and Raginsky (2022); Kotelevskii et al.
+(2025).
+Recently, this connection was revisited in the context of natural language generation Wang and Holmes (2025); Daheim et al.
+(2025), leading to the development of new MBR-based UQ methods.
+More precisely, one can consider the smallest achievable risk within $\mathcal{Y}$ or, equivalently, to maximum expected utility:
+$U_{\text{MBR}}(\mathbf{y}_{*}\mid\mathbf{x})=\min_{\mathbf{y}\in\mathcal{Y}}R(\mathbf{y}\mid\mathbf{x})=R(\mathbf{y}_{*}\mid\mathbf{x}),$ (4)
+where $\mathbf{y}_{*}$ is given by (2). Interestingly, MBR-based uncertainty (4) is applicable even to $\mathcal{Y}$ consisting of a single sequence $\mathbf{y}_{*}$.
+Below, we discuss how various existing uncertainty measures can be seen as particular approximations of minimum Bayes risks for various choices of utility functions.
+We also derive new uncertainty measures based on the MBR framework.
+Single Sequence Information-Based Methods.
+Information-based methods estimate the uncertainty of the generated sequence by aggregating the uncertainty scores of individual tokens.
+Within the framework of MBR decoding, consider $r(\mathbf{y},\mathbf{y}^{\prime})=\mathbf{1}\{\mathbf{y}^{\prime}\neq\mathbf{y}\}$.
+In this case, Bayes risk corresponds to the expected zero-one loss of decoding:
+$R_{0/1}(\mathbf{y}\mid\mathbf{x})=\mathbb{E}_{\mathbf{y}^{\prime}\sim p(\mathbf{y}\mid\mathbf{x})}\mathbf{1}\{\mathbf{y}^{\prime}\neq\mathbf{y}\}=1-p(\mathbf{y}\mid\mathbf{x}).$ (5)
+This leads to one of the simplest information-based uncertainty measures, Maximum Sequence Probability (MSP):
+$U_{\text{MSP}}(\mathbf{y}_{*}\mid\mathbf{x})=1-p\bigl{(}\mathbf{y}_{*}\mid\mathbf{x}\bigr{)}.$ (6)
+Although widely used as an uncertainty measure across various applications, including natural language generation, its connection to statistical risk has only recently been explored in the UQ literature Kotelevskii et al.
+(2022, 2025); Aichberger et al.
+(2024).
+Several other measures fall into this category, including Perplexity and Mean Token Entropy (Fomicheva et al., 2020); see Appendix D.1 for details.
 While using only a single sample makes them computationally efficient, these techniques face three major challenges:
   1. 1.
-    First, LLMs provide us with the probability of a specific answer, even though the same meaning could often be expressed in multiple ways.
-Therefore, to obtain a proper probability for the meaning of an answer, we need to marginalize over its various possible rephrasings.
+    LLMs provide us only the probability of a specific answer, even though the same meaning can often be conveyed in multiple ways.
+To obtain a proper probability for the meaning of an answer, we need to marginalize over its various possible rephrasings.
 However, this is not feasible if we generate only a single sample.
   2. 2.
-    Second, answers of LLMs in general have very low probabilities due to the size of the vocabulary and the length of the generated sequences.
-Therefore, their estimates by LLM are not very reliable in general.
+    Reliability of such methods depends heavily on the underlying LLM being well-calibrated, which is a quality that is hard to define and harder to ensure.
   3. 3.
-    Third, these methods do not provide information about the flatness or the variability of the answer distribution.
-2.2 Consistency-Based Methods
+    These methods are point estimates that do not provide information about the shape of the output distribution.
+Semantic Consistency-Based Methods.
 The aforementioned issues lead to the development of consistency-based methods based on repetitive sampling from the LLM.
 Consider that we have sampled a set of outputs $\bigl{\{}\mathbf{y}^{(i)}\bigr{\}}_{i=1}^{M}$, where ${\mathbf{y}^{(i)}\sim p(\mathbf{y}\mid\mathbf{x})}$.
-Consistency-based uncertainty quantification methods rely only on the diversity of answers $\mathbf{y}^{(i)}$ sampled from the LLM.
+Consistency-based UQ methods rely on the diversity of answers $\mathbf{y}^{(i)}$ sampled from the LLM.
 The idea is that if the model outputs similar answers for the same prompt over and over again, it is confident in its predictions; otherwise, it is uncertain.
-These techniques do not require knowledge about the probability distribution of the tokens and can be applied in the black-box setting, when only the generated tokens are available.
+These techniques do not require a probability distribution estimated by LLM and can be applied in the black-box setting, when only the generated tokens are available.
 This case is quite common when LLMs are deployed as a service and are accessible through a limited API.
-Formally, given $M$ samples from the model, consistency-based methods compute a similarity matrix $G$, where each element $g_{ij}$ represents some form of similarity between the sampled outputs $\mathbf{y}^{(i)}$ and $\mathbf{y}^{(j)}$:
-$g\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}\in[0,1].$ (3)
-The value $g\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}=1$ indicates the complete equivalence between $\mathbf{y}^{(i)}$ and $\mathbf{y}^{(j)}$, and $g\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}=0$ indicates that there is no similarity.
+Formally, consistency-based method start from defining some similarity function $s\bigl{(}\mathbf{y},\mathbf{y}^{\prime}\bigr{)}\in[0,1]$ between arbitrary generations $\mathbf{y}$ and $\mathbf{y}^{\prime}$.
+The value $s\bigl{(}\mathbf{y},\mathbf{y}^{\prime}\bigr{)}=1$ indicates the complete equivalence between $\mathbf{y}$ and $\mathbf{y}^{\prime}$, and $s\bigl{(}\mathbf{y},\mathbf{y}^{\prime}\bigr{)}=0$ indicates that there is no similarity.
 Similarity could be computed in various ways.
-For example, Lexical Similarity (Fomicheva et al., 2020) is the surface form similarity, which calculates the overlap of words or phrases in the generations.
+For instance, the Lexical Similarity method Fomicheva et al.
+(2020) relies on surface-form similarity, measuring the degree of word- or phrase-level overlap between generated texts.
 More advanced techniques propose various methods for taking into account the semantic similarity of the generated answers by hard or soft clustering (Lin et al., 2024).
-For example, the Degree Matrix approach considers a similarity matrix $G$, which is computed using a model for Natural Language Inference (NLI), which predicts the probabilities of entailment, ${p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})$, and contradiction, ${p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$, between pairs of sentences $\mathbf{y}$ and $\mathbf{y}^{\prime}$.
-The similarity between two sequences is then defined as either $g_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})={p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})\quad\text{or}\quad g_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})=1-{p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$.
-Lin et al.
-(2024) proposes to compute an averaged similarity matrix as $g_{ij}=\bigl{(}g\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}+g\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(i)}\bigr{)}\bigr{)}/2$.
-The diagonal matrix $D$ is then defined with elements $D_{ii}=\sum_{j=1}^{M}g_{ij}$.
-The uncertainty from this soft clustering of meanings can then be defined as follows:
-$U_{\mathrm{DegMat}}=1-\frac{\mathrm{trace}(D)}{M^{2}}.$ (4)
-The advantage of these techniques is that by generating multiple samples and analyzing their semantic similarity, they can obtain empirical probabilities for meanings instead of individual answers.
+Finally, given $M$ samples from the model, standard consistency-based methods compute a similarity matrix $S$, where $s_{ij}=s\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}$.
+Then, various statistics of $S$ are computed in order to estimate uncertainty (Lin et al., 2024), see details in Appendix D.2.
+Alternatively, one can focus on a particular output $\mathbf{y}_{*}$ and compute uncertainty measures based on its similarities $s_{*i}=s\bigl{(}\mathbf{y}_{*},\mathbf{y}^{(i)}\bigr{)}$ with other samples $\mathbf{y}^{(i)}$; see Lin et al.
+(2024) for more details.
+Interestingly, it was recently shown Wang and Holmes (2025); Daheim et al.
+(2025) that the latter approach has direct relation to MBR decoding.
+Building on the MBR framework, one can define the utility function as $r(\mathbf{y},\mathbf{y}^{\prime})=1-s(\mathbf{y},\mathbf{y}^{\prime})$ and derive the following consistency-based uncertainty score:
+$U_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})=\mathbb{E}_{\mathbf{y}^{\prime}\sim p(\mathbf{y}\mid\mathbf{x})}\bigl{[}1-s(\mathbf{y}_{*},\mathbf{y}^{\prime})\bigr{]}.$ (7)
+Then the Monte Carlo approximation of the minimum Bayes risk is given by:
+$\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})=\frac{1}{M}\sum\nolimits_{i=1}^{M}(1-s_{*i}),$ (8)
+where $s_{*i}=s\bigl{(}\mathbf{y}_{*},\mathbf{y}^{(i)}\bigr{)}$.
+In our ablation study, we show that such an uncertainty measure reliably outperforms consistency-based measures that aggregate the pairwise similarities of all the samples (see Appendix C.2).
+A key strength of consistency-based techniques is that, by generating multiple samples and analyzing their semantic similarity, they can estimate empirical probabilities over meanings rather than individual answers.
 The main drawback is that they discard the useful information that comes from the probability distribution represented by the LLM, including estimates of probabilities of specific answers.
-2.3 Information-Based Methods with Repeated Sampling
-The natural idea is to somehow benefit from having multiple samples from the model while using important information contained in the output probabilities estimated by an LLM.
-Below, we examine several approaches that have sought to achieve this.
-Averaging uncertainties.
-The uncertainty scores can be aggregated using simple Monte Carlo averaging:
-$u_{\text{MC}}=\frac{1}{M}\sum_{i=1}^{M}u_{i}.$ (5)
-For the case when using the MSP uncertainty measure, i.e., when $u_{i}=-\log p\bigl{(}\mathbf{y}^{(i)}\mid\mathbf{x}\bigr{)}$, we obtain $u_{\text{MC}}=-\frac{1}{M}\sum_{i=1}^{M}\log p\bigl{(}\mathbf{y}^{(i)}\mid\mathbf{x}\bigr{)}$.
-The other notable example is the Monte Carlo Sequence Entropy (Kuhn et al., 2023).
-While simple averaging represents a natural way to aggregate uncertainties, it has certain issues related to the nature of LLMs.
-First of all, in the vast majority of applications, an LLM-based system should produce a single output $\mathbf{y}_{*}$ for an input query.
-When we consider $u_{\text{MC}}$, we essentially perform averaging of uncertainties of different sequences, thus somewhat assessing the uncertainty related to the entire generative distribution $p(\mathbf{y}\mid\mathbf{x})$ for the input $\mathbf{x}$, but not for a particular generated sequence $\mathbf{y}_{*}$.
-This averaged uncertainty might not be adequate for this particular sequence and, remarkably, often performs worse than the uncertainty $u_{*}=U(\mathbf{y}_{*}\mid\mathbf{x})$, which is related solely to the output $\mathbf{y}_{*}$.
-Moreover, although intuitive, this naïve aggregation method assumes that all outputs contribute equally to the final uncertainty estimate, regardless of their semantic relationships.
-This can lead to inconsistencies when semantically equivalent outputs have varying uncertainty scores or when outputs with low similarity are treated as equally important.
-Semantically weighted averaging.
-The basic idea of aggregation approaches like Semantic Entropy (Kuhn et al., 2023) or SAR (Duan et al., 2024) is to perform a weighted averaging of output probabilities and give more weight to sequences semantically similar to the response shown to a user.
-All recently proposed techniques, such as SAR and Semantic Entropy, can be unified into a semantically-aware Generalized Monte Carlo uncertainty estimate, defined as
-$u_{\text{GMCU}}=\frac{1}{M}\sum_{i=1}^{M}h\Biggl{(}\sum_{j=1}^{M}g_{ij}\,p_{j}\Biggr{)}.$ (6)
-Here, the inner summation aggregates sequence probabilities $p_{j}$ weighted by their semantic similarity to the $i$-th output, and the outer summation averages these contributions across all samples.
-The function $h(\cdot)$ provides an additional layer of flexibility, transforming the reweighted uncertainty scores, making the method a generalized framework for uncertainty quantification.
-Existing methods, such as Semantic Entropy and SAR, can be considered as special cases of this more comprehensive approach, where the functions $h$ and $g$ are chosen appropriately.
-Unfortunately, methods that fall under GMCU, while offering benefits, also inherit the aforementioned issues from both categories of methods:
-  1. 1.
-    The term $\sum_{j=1}^{M}g_{ij}\,p_{j}$ aims to average the probabilities of semantically similar sequences to obtain more robust estimate of the probability.
-However, due to the extreme instability of the LLM probabilities, as shown in Figure 1, the aggregated probabilities often perform worse than non-aggregated baselines.
-  2. 2.
-    The outer summation in (6), similarly to the case of simple Monte Carlo averaging (5), often fails to outperform the uncertainty $u_{*}=U(\mathbf{y}_{*}\mid\mathbf{x})$ of a single generated sequence $\mathbf{y}_{*}$.
-Let us note that all uncertainty functions discussed so far have the following properties:
-  • Non-Negativity. The uncertainty function produces nonnegative values, that is, $U(\mathbf{y})\geq 0$ for all $\mathbf{y}$.
-  • Monotonicity.
-Higher values of the uncertainty function $U$ indicate higher uncertainty, i.e., if output $\mathbf{y}^{(1)}$ is considered more uncertain than output $\mathbf{y}^{(2)}$, then $U\bigl{(}\mathbf{y}^{(1)}\bigr{)}\geq U\bigl{(}\mathbf{y}^{(2)}\bigr{)}$.
-These properties become important to perform the kind of synthesis of confidence and consistency we propose in the following section.
+Figure 2: Illustration of the method: the LLM generates a response, evaluates the similarity to alternatives, computes confidence, and combines confidence with the similarity measure.
+High similarity to alternatives reduces the uncertainty, while low similarity keeps it high.
 3 CoCoA: Bridging Confidence and Consistency for Better Uncertainty Quantification
-We start by summarizing the benefits and drawbacks of various uncertainty quantification approaches discussed above:
-  1. 1.
 Both information-based and (semantic) consistency-based methods provide grounded and useful uncertainty quantification measures.
-  2. 2.
-    Output probabilities $p(\mathbf{y}^{(j)}\mid\mathbf{x}),\,j=1,\dots,M$ might have substantially different values for semantically equivalent outputs, which questions the usefulness of (weighted) averaging these probabilities for uncertainty quantification.
-  3. 3.
-    For various methods based on the aggregation over multiple samples, the result might be suboptimal due to the noise related to the averaging over all generated outputs.
-Focusing solely on a particular output sequence and its relation to other generated outputs might be beneficial.
-In what follows, we present a family of UQ Confidence and Consistency-based Approaches (CoCoA), offering a new way to merge information- and consistency-based measures for uncertainty quantification in LLMs.
-Let us consider an actual output sequence $\mathbf{y}_{*}$ and a set of sampled sequences $\mathbf{y}^{(i)},\,i=1,\dots,M$.
-Here, $\mathbf{y}_{*}$ might be one of the sequences $\mathbf{y}^{(i)}$ or might be generated separately.
-In what follows, we will consider several possible cases, including $\mathbf{y}_{*}$ being a random sequence from a set $\{\mathbf{y}^{(i)}\}$, $\mathbf{y}_{*}$ being a sequence from a set $\{\mathbf{y}^{(i)}\}$ having the highest probability, and, finally, $\mathbf{y}_{*}$ being a sequence found via the beam search procedure.
-First, consider an information-based uncertainty score of the output $\mathbf{y}_{*}$:
-$u_{*}^{\text{info}}=U^{\text{info}}(\mathbf{y}_{*}\mid\mathbf{x}),$ (7)
-where $U^{\text{info}}$ might be MSP, perplexity, mean token entropy, or another uncertainty measure related solely to the generated sequence $\mathbf{y}_{*}$.
-We quantify the consistency-based uncertainty via a direct measurement of the semantic similarity of generated sequence $\mathbf{y}_{*}$ to sampled sequences:
-$u_{*}^{\text{cons}}=\frac{1}{M}\sum_{i=1}^{M}(1-g_{*i}),$ (8)
-where $g_{*i}=g\bigl{(}\mathbf{y}_{*},\mathbf{y}^{(i)}\bigr{)}$.
-This formulation satisfies the desired properties of the uncertainty function – that is their values are nonnegative and their values increase with increased inconsistency (decreasing value of $g_{*i}$).
-In our ablation study, we will show that such an uncertainty measure reliably outperforms consistency-based measures that aggregate the pairwise similarities of all the samples (see Appendix B.3).
-Finally, we need to aggregate $u_{*}^{\text{info}}$ and $u_{*}^{\text{cons}}$ into a single uncertainty measure. We propose to aggregate them in a multiplicative way:
-$u_{*}^{\texttt{CoCoA}}=u_{*}^{\text{info}}\cdot u_{*}^{\text{cons}}$ (9)
-This formulation preserves the non-negativity and the monotonicity properties while integrating both global (semantic) and local (model-specific) uncertainty signals.
-It ensures that uncertainty is amplified for sequences that are both intrinsically uncertain (high $u_{*}^{\text{info}}$) and semantically inconsistent with the dataset (high $u_{*}^{\text{cons}}$), while keeping it low for the opposite scenario (see Figure 2).
-Although the choice of the multiplicative aggregation function CoCoA is heuristic, it provides a practical and effective way to combine information- and consistency-based uncertainty signals in LLMs.
-In our ablation study, we also compare the multiplicative formulation in (9) to a simpler additive variant, $u_{*}^{\text{info}}+u_{*}^{\text{cons}}$ (see Appendix B.3).
-Empirically, the multiplicative combination is better at capturing the joint impact of both information-based and consistency-based uncertainty, yielding more reliable estimation across all tasks.
+There exist approaches that bridge the gap between information-based and consistency-based methods that show great promise but lack the fundamental base; see discussion in Section 4.
+In what follows, we present a family of Confidence and Consistency-based Approaches (CoCoA) for UQ, offering a new way to merge information- and consistency-based measures for uncertainty quantification in LLMs via the unified MBR-based framework.
+If we consider only the consistency measure, like the one in (8), we omit the information that is contained in the model confidence.
+Thus, we want the resulting uncertainty measure to explicitly consider both semantic consistency and model-based uncertainty.
+Let us consider the risk of the form:
+$r(\mathbf{y},\mathbf{y}^{\prime}\mid\mathbf{x})=u(\mathbf{y}\mid\mathbf{x})\cdot\bigl{(}1-s(\mathbf{y},\mathbf{y}^{\prime})\bigr{)},$ (9)
+where
+  • $s(\mathbf{y},\mathbf{y}^{\prime})\in[0,1]$ is any utility function used in the standard MBR decoding (usually representing semantic similarity);
+  • $u(\mathbf{y}\mid\mathbf{x})\geq 0$ is the model-based uncertainty measure for the output sequence $\mathbf{y}$.
+For example, we can use $u(\mathbf{y}\mid\mathbf{x})=1-p(\mathbf{y}\mid\mathbf{x})$ (similarly to (6)) or $u(\mathbf{y}\mid\mathbf{x})=-\log p(\mathbf{y}\mid\mathbf{x})$.
+Then, the corresponding Bayes risk is $R_{\text{CoCoA}}(\mathbf{y}\mid\mathbf{x})=u(\mathbf{y}\mid\mathbf{x})\cdot\mathbb{E}_{\mathbf{y}^{\prime}\sim p(\mathbf{y}\mid\mathbf{x})}\,\bigl{(}1-s(\mathbf{y},\mathbf{y}^{\prime})\bigr{)}$ and for an output sequence $\mathbf{y}_{*}$, the resulting uncertainty measure becomes:
+$U_{\text{CoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})\cdot\mathbb{E}_{\mathbf{y}^{\prime}\sim p(\mathbf{y}\mid\mathbf{x})}\,\bigl{(}1-s(\mathbf{y}_{*},\mathbf{y}^{\prime})\bigr{)}.$ (10)
+Finally, given a set of samples $\bigl{\{}\mathbf{y}^{(i)}\bigr{\}}_{i=1}^{M}$ we obtain an empirical uncertainty estimate:
+$\widehat{U}_{\text{CoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})\cdot\frac{1}{M}\sum\nolimits_{i=1}^{M}(1-s_{*i})=u(\mathbf{y}_{*}\mid\mathbf{x})\cdot\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x}),$ (11)
+where $s_{*i}=s\bigl{(}\mathbf{y}_{*},\mathbf{y}^{(i)}\bigr{)}$.
+The resulting uncertainty measure integrates both global (semantic) and local (model-specific) uncertainty signals.
+It ensures that uncertainty is amplified for sequences $\mathbf{y}_{*}$ that are both intrinsically uncertain (high $u(\mathbf{y}_{*}\mid\mathbf{x})$) and semantically inconsistent with the other samples (high $\frac{1}{M}\sum\nolimits_{i=1}^{M}(1-s_{*i})$), while keeping it low for the opposite scenario; see Figure 2.
+CoCoA Light.
+Computing the consistency-based uncertainty measure $\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})$ within $\widehat{U}_{\text{CoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})$ requires multiple samples, leading to significant computational overhead.
+To address this, we propose to approximate the behavior of $\widehat{U}_{\text{cons}}(\mathbf{x})$ with a learned function.
+Our method closely matches the original sampling-based measure in quality while requiring only greedy output generation during inference, eliminating the need for additional sampling.
+Notably, this approximation can be learned without access to ground-truth labels, relying solely on the input data and the associated uncertainty values.
+Thus, the CoCoA uncertainty measure that incorporates an approximation of the consistency uncertainty $\widehat{U}^{L}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})\approx\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})$ can be defined as:
+$\widehat{U}^{L}_{\text{CoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})\cdot\widehat{U}^{L}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x}),$ (12)
+We name this approach CoCoA Light. Below, we describe how this approximation can be obtained.
+Assume access to a held-out set of input sequences $\mathbf{x}_{j},j=1,\dots,n$.
+We emphasize that this held-out set is not labeled.
+For each sequence, we extract the embeddings $\{e(\mathbf{y}_{j}\mid\mathbf{x}_{j})\}_{j=1}^{n}$ of the corresponding greedy model generations $\mathbf{y}_{j}$.
+The corresponding targets are the consistency uncertainty scores ${\widehat{U}_{\text{cons}}(\mathbf{y}_{j}\mid\mathbf{x}_{j})},j=1,\dots,n$, computed via Monte Carlo estimation of the minimum Bayes risk; see equation (8).
+A lightweight auxiliary model, $\widehat{g}\bigl{(}e(\mathbf{y}_{*}\mid\mathbf{x})\bigr{)}$, is trained in a supervised fashion to map model embeddings to uncertainty scores.
+During inference, the model generates a greedy output $\mathbf{y}_{\text{test}}$ for a test input $\mathbf{x}_{\text{test}}$, from which the embeddings $e(\mathbf{y}_{\text{test}}\mid\mathbf{x}_{\text{test}})$ are obtained and passed to the auxiliary model to predict the uncertainty score:
+$\widehat{U}^{L}_{\text{cons}}(\mathbf{y}_{\text{test}}\mid\mathbf{x}_{\text{test}})=\widehat{g}\bigl{(}e(\mathbf{y}_{\text{test}}\mid\mathbf{x}_{\text{test}})\bigr{)}.$ (13)
+We note that a similar approach (Kossen et al., 2024) has been previously proposed to approximate semantic entropy using hidden states from the model.
+4 Related Work
+Information-based methods.
+These methods are one of the most commonly used.
+They quantify uncertainty by analyzing the probability distributions of tokens within a given output.
+The simplest of these methods, Maximum Sequence Probability (MSP), measures the probability of the most likely output sequence given a specific input.
+Perplexity is another common uncertainty measure: it gauges how well the model predicts each token in a sequence and is formally defined as the exponential of the mean negative log-likelihood of those tokens (Fomicheva et al., 2020).
+The Mean Token Entropy method evaluates token-level predictions across the entire sequence by computing the average entropy of the token probability distributions at each position (Fomicheva et al., 2020).
+While these methods provide useful uncertainty estimates, they do not address the broader uncertainty inherent in generative tasks, as they rely on a single sequence and fail to capture the diversity of possible outputs that the model could generate for the same input.
+Consistency-Based Methods.
+These methods estimate uncertainty by generating multiple sequences from the model’s output distribution and analyzing the variability among the sampled outputs.
+These methods are particularly relevant to NLP tasks, as they account for scenarios in which multiple plausible outputs may exist for a given input.
+Among the simplest consistency-based methods are the Number of Semantic Sets and the Sum of Eigenvalues of the Graph Laplacian Lin et al.
+(2024), which quantify uncertainty by how many distinct “meanings” the model produces.
+While effective at capturing global variation, they do not yield uncertainty scores for individual responses.
+To overcome this, the diagonal of the Degree Matrix was proposed to assess how similar each output is to the rest, enabling per-response uncertainty estimation Lin et al.
+(2024).
+A related approach, Lexical Similarity, operates by calculating the average similarity of words or phrases across every pair of responses in the sample Fomicheva et al.
+(2020).
+Information-Based Methods with Repeated Sampling.
+More recent methods attempt to reconcile output consistency with information-based signals—i.e., they not only look at how varied outputs are, but also how probable each output is under the model’s own generation process.
+For instance, Semantic Entropy (Kuhn et al., 2023) groups outputs into semantically homogeneous clusters—capturing distinct “meanings” that might just differ in phrasing—and calculates entropy across these clusters.
+SentenceSAR refines this idea by defining “relevance” as the sum of pairwise similarities between a given sentence and others weighted by each sentence’s model-assigned generation probability (Duan et al., 2024).
+SAR combines the sentence-level relevance of SentenceSAR with token-level probability adjustments, attempting a more fine-grained balance of semantic and token-level uncertainty (Duan et al., 2024).
+Semantic Density, Qiu and Miikkulainen (2024) evaluates uncertainty by evaluating how densely a model’s generated response is situated within the semantic space of all possible outputs, with lower density indicating higher uncertainty.
+One limitation of sampling-based approaches is their computational cost: generating multiple outputs from the model can be expensive.
+Recent work has sought to address this by learning to predict semantic uncertainty directly from hidden states, allowing for the approximation of semantic entropy without repeated sampling Kossen et al.
+(2024).
+Finally, BSDetector (Chen and Mueller, 2024) proposes an additive composition of self-reported confidence with observed consistency, however it relies on questionable ability of the model to assess it’s own confidence in text form.
+Another limitation is that it requires selection of a trade-off coefficient between confidence and consistency parts.
+While these probability-weighted measures can provide deeper insights, they also sometimes underperform in practice (Vashurin et al., 2025), especially when the model’s probability estimates are unreliable, or when important nuances of semantic diversity are lost in the weighting process.
+Proper balancing raw output consistency with generation likelihood remains an open problem.
+Minimum Bayes Risk for LLMs.
+Recent work has proposed incorporating uncertainty into Minimum Bayes Risk decoding by introducing a posterior over model parameters, enabling uncertainty-aware generation Daheim et al.
+(2025).
+Complementing this, a Bayesian framework for subjective uncertainty quantification was developed yielding improved calibration in generation tasks (Wang and Holmes, 2025).
 5 Experiments
 5.1 Experimental Setup
-To evaluate the effectiveness of the proposed method, we extended the LM-Polygraph library (Vashurin et al., 2024; Fadeeva et al., 2023) by implementing our approach within its framework.
+To evaluate the effectiveness of the proposed method, we extended the LM-Polygraph library (Vashurin et al., 2025; Fadeeva et al., 2023) by implementing our approach within its framework.
 Since the library already includes tools for calculating other uncertainty scores, it provided a convenient and efficient environment for setting up and running experiments.
 The primary objective of our experiments is to evaluate whether our method offers improved performance in key tasks such as question answering (QA), summarization (SUM), and machine translation (MT), compared to existing baselines.
-Figure 3: Prediction-Rejection Ratio (PRR) Curve.
-The curve illustrates the quality of the non-rejected predictions as a function of the rejection rate.
-Oracle represents the optimal rejection strategy, Random is a random rejection, and UE is rejection based on the evaluated uncertainty estimation method.
 Datasets.
@@
 Finally, for summarization, we used XSUM (Narayan et al., 2018), a dataset of complex documents paired with concise abstractive summaries.
-For all datasets, we follow (Vashurin et al., 2024) for selecting the subsets, for prompt formatting, and for number and sourcing of few-shot examples.
+For all datasets, we follow (Vashurin et al., 2025) for selecting the subsets, for prompt formatting, and for the number and sourcing of few-shot examples.
+Figure 3: Prediction-Rejection Ratio (PRR) Curve.
+The curve illustrates the quality of the non-rejected predictions as a function of the rejection rate.
+Oracle represents the optimal rejection strategy, Random is a random rejection, and UE is rejection based on the evaluated uncertainty estimation method.
 Models.
-We evaluated our method using the base versions of three open-weights language models: LLaMA 3.1 8B (Touvron et al., 2023), Mistral 7B (Jiang et al., 2023), and Falcon 3 7B (Team, 2024).
+We evaluated our method using the base versions of three open-weight language models: LLaMA 3.1 8B Touvron et al.
+(2023), Mistral 7B Jiang et al.
+(2023), and Falcon 3 7B Team (2024).
 The open-source nature of these models provides full access to their token probabilities, which are essential for implementing our UQ method.
-For all models, we consider base versions, without instruction-tuning.
-Metric Llama Mistral Falcon
-QA NMT SUM QA NMT SUM QA NMT SUM
-MCSE 0.357 0.380 0.192 0.453 0.406 0.162 0.460 0.409 0.128
-MCNSE 0.380 0.429 0.186 0.466 0.489 0.196 0.530 0.424 0.153
-Semantic Entropy 0.397 0.411 0.194 0.482 0.438 0.164 0.479 0.440 0.134
-SAR 0.479 0.506 0.159 0.542 0.576 0.175 0.590 0.488 0.193
-DegMat 0.422 0.342 0.191 0.465 0.425 0.205 0.543 0.386 0.177
-EigValLaplacian 0.388 0.274 0.190 0.426 0.366 0.197 0.498 0.336 0.174
-MSP 0.395 0.376 0.464 0.444 0.252 0.330 0.343 0.381 0.099
-$\text{CoCoA}_{MSP}$ 0.484 $\uparrow$ 0.607 $\uparrow$ 0.484 $\uparrow$ 0.526 $\uparrow$ 0.721 $\uparrow$ 0.366 $\uparrow$ 0.529 $\uparrow$ 0.631 $\uparrow$ 0.210 $\uparrow$
-PPL 0.532 0.563 0.458 0.587 0.686 0.365 0.627 0.589 0.275
-$\text{CoCoA}_{PPL}$ 0.571 $\uparrow$ 0.617 $\uparrow$ 0.450 0.613 $\uparrow$ 0.745 $\uparrow$ 0.372 $\uparrow$ 0.647 $\uparrow$ 0.648 $\uparrow$ 0.310 $\uparrow$
-NMTE 0.477 0.469 0.449 0.559 0.637 0.350 0.602 0.492 0.186
-$\text{CoCoA}_{NMTE}$ 0.548 $\uparrow$ 0.579 $\uparrow$ 0.451 $\uparrow$ 0.600 $\uparrow$ 0.720 $\uparrow$ 0.373 $\uparrow$ 0.641 $\uparrow$ 0.614 $\uparrow$ 0.289 $\uparrow$
-Table 1: Results for Evaluated Sequence – Best Sample: Mean PRR across datasets for each task.
-The best performing method is in bold, and the second-best is underscored.
-Arrows indicate improvement in CoCoA over the base version.
+For all models, we consider base versions, without instruction tuning.
 Similarity Function.
@@
 The cross-encoder processes two sequences jointly and directly outputs a similarity score ranging from 0 to 1, providing a nuanced measure.
-Appendix B.2 contains comparative experiments with cross-encoder and other choices of the similarity function, substantiating this choice.
+Appendix C.1 contains comparative experiments with cross-encoder and other choices of the similarity function, substantiating this choice.
 Baselines.
-We compare the performance of the proposed method against a diverse set of baseline and state-of-the-art UQ scores, including confidence-based, consistency-based, and hybrid approaches.
+We compare the performance of the proposed method against a diverse set of baselines and state-of-the-art UQ scores, including confidence-based, consistency-based, and hybrid approaches.
 Information-based methods include Maximum Sequence Probability (MSP), Perplexity (PPL), Mean Token Entropy (MTE), Monte Carlo Sequence Entropy (MCSE), and Monte Carlo Normalized Sequence Entropy (MCNSE).
@@
 Finally, hybrid methods include Semantic Entropy and SAR.
-All formulations for these baselines can be found in Appendix C.
+All formulations for these baselines can be found in Appendix D.
+Lastly, we evaluate performance of $\hat{U}_{\text{cons}}$ (Consistency) and $\hat{U}_{\text{cons}}^{L}$ (Consitency Light) as an uncertainty measure.
 Evaluation measure.
 As our evaluation measure, we choose the Prediction Rejection Ratio (PRR), which measures the effectiveness of the uncertainty scores for identifying high-quality predictions (Malinin and Gales, 2021).
-PRR operates by progressively rejecting predictions with uncertainty scores above a threshold $a$ and observing how the average quality $Q(f(\mathbf{x}_{i}),\mathbf{y}_{i})$ of the remaining predictions changes.
+PRR operates by progressively rejecting predictions with uncertainty scores above a threshold $a$ and observing how the average quality $Q(f(\mathbf{x}_{i}),\mathbf{y}_{i})$ of the remaining predictions changes (Figure 3).
 The metric is calculated as the ratio of two areas: the area between the Prediction Rejection (PR) curves for the evaluated uncertainty score and a random baseline, and the area between the oracle (the ideal uncertainty score that perfectly ranks instances by quality) and the random baseline.
@@
 To ensure practical applicability, we compute PRR only up to a rejection threshold of 50%, preventing cases where excessive rejection artificially inflates quality measures.
-Figure 3 gives a visual representation of the PRR calculation, highlighting the relationship between the uncertainty threshold and the quality measures.
-Metric Llama Mistral Falcon
+For some of the datasets we also report AUROC results, see Appendix G.
+Quality Measures.
+PRR requires an appropriate quality measure for each specific task to effectively evaluate the model output.
+For question-answering tasks, we use Accuracy to directly evaluate whether the generated answers match the ground truth in short-form QA tasks (e.g., MMLU), and we use the AlignScore between the correct answer and generated sequence for assessing the performance for long-form QA tasks (Zha et al., 2023).
+For summarization tasks, we use AlignScore to measure the alignment between the output summary and the input document.
+It serves as a quality indicator by evaluating the relevance and the overlap between the generated content and the source text.
+For translation tasks, we use COMET, as it captures both semantic adequacy and fluency, ensuring that translations are accurate and linguistically appropriate (Rei et al., 2020).
+To further improve comprehensiveness of reported results, we considered alternative choices of quality measures, and report corresponding PRR scores in Appendix F.
+Generation Setup.
+We discuss the generation parameters, the decoding strategy, and sample selection procedure in depth in Appendix A.
+In short, we report evaluation results in two distinct setups: greedy decoding and stochastic sampling with a focus on the most probable sequence among the generated outputs (best-sample).
+These two setups offer the highest-quality outputs and are the most reasonable generation approaches in practice.
+CoCoA Light Details.
+The multilayer perceptron is used as an auxiliary model.
+The features used are embeddings from the middle layers of the base LLM, which are the most informative according to recent work Chen et al.
+(2024).
+In particular, for Llama 3.1 8B and Mistral 7B models, embeddings are taken from the 16th layer, and for Falcon 3 7B from the 14th layer.
+Metric Llama8b-Base Mistral7b-Base Falcon7b-Base
 QA NMT SUM QA NMT SUM QA NMT SUM
@@
 Semantic Entropy 0.356 0.343 0.033 0.423 0.327 0.008 0.439 0.348 0.164
-SAR 0.414 0.455 0.077 0.462 0.435 0.094 0.481 0.458 0.144
 DegMat 0.406 0.302 0.081 0.423 0.305 0.137 0.483 0.353 0.201
 EigValLaplacian 0.375 0.238 0.079 0.391 0.267 0.132 0.459 0.312 0.201
+SAR 0.414 0.455 0.077 0.462 0.435 0.094 0.481 0.458 0.144
+Consistency 0.437 0.421 0.024 0.471 0.392 0.051 0.494 0.416 0.226
+Consistency Light 0.390 0.458 -0.022 0.444 0.387 -0.006 0.427 0.476 0.232
 MSP 0.409 0.399 0.328 0.475 0.383 0.287 0.475 0.356 0.201
 $\text{CoCoA}_{MSP}$ 0.451 $\uparrow$ 0.519 $\uparrow$ 0.378 $\uparrow$ 0.509 $\uparrow$ 0.497 $\uparrow$ 0.330 $\uparrow$ 0.511 $\uparrow$ 0.505 $\uparrow$ 0.257 $\uparrow$
+$\text{CoCoA}_{MSP}\,\,\text{Light}$ 0.449 $\uparrow$ 0.502 $\uparrow$ 0.358 $\uparrow$ 0.503 $\uparrow$ 0.480 $\uparrow$ 0.309 $\uparrow$ 0.492 $\uparrow$ 0.514 $\uparrow$ 0.242 $\uparrow$
 PPL 0.381 0.386 0.369 0.424 0.427 0.204 0.456 0.450 0.155
 $\text{CoCoA}_{PPL}$ 0.454 $\uparrow$ 0.481 $\uparrow$ 0.387 $\uparrow$ 0.494 $\uparrow$ 0.472 $\uparrow$ 0.286 $\uparrow$ 0.523 $\uparrow$ 0.508 $\uparrow$ 0.229 $\uparrow$
-NMTE 0.353 0.382 0.357 0.417 0.438 0.182 0.456 0.473 0.152
-$\text{CoCoA}_{NMTE}$ 0.447 $\uparrow$ 0.478 $\uparrow$ 0.380 $\uparrow$ 0.492 $\uparrow$ 0.469 $\uparrow$ 0.288 $\uparrow$ 0.527 $\uparrow$ 0.508 $\uparrow$ 0.228 $\uparrow$
+$\text{CoCoA}_{PPL}\,\,\text{Light}$ 0.445 $\uparrow$ 0.487 $\uparrow$ 0.382 $\uparrow$ 0.479 $\uparrow$ 0.480 $\uparrow$ 0.260 $\uparrow$ 0.512 $\uparrow$ 0.528 $\uparrow$ 0.234 $\uparrow$
+MTE 0.353 0.382 0.357 0.417 0.438 0.182 0.456 0.473 0.152
+$\text{CoCoA}_{MTE}$ 0.447 $\uparrow$ 0.478 $\uparrow$ 0.380 $\uparrow$ 0.492 $\uparrow$ 0.469 $\uparrow$ 0.288 $\uparrow$ 0.527 $\uparrow$ 0.508 $\uparrow$ 0.228 $\uparrow$
+$\text{CoCoA}_{MTE}\,\,\text{Light}$ 0.428 $\uparrow$ 0.494 $\uparrow$ 0.372 $\uparrow$ 0.475 $\uparrow$ 0.482 $\uparrow$ 0.254 $\uparrow$ 0.507 $\uparrow$ 0.533 $\uparrow$ 0.234 $\uparrow$
 Table 1: Results for Evaluated Sequence – Greedy Sample: Mean PRR across datasets for each task.
@@
 Arrows indicate improvement in CoCoA over the base version.
-Quality Measures.
-The Predictive Rejection Ratio (PRR) requires an appropriate quality measure for each specific task to effectively evaluate the model output.
-For question-answering tasks, we use Accuracy to directly evaluate whether the generated answers match the ground truth in short-form QA tasks (e.g., MMLU), and we use the AlignScore between correct answer and generated sequence for assessing the performance for long-form QA tasks (Zha et al., 2023).
-For summarization tasks, we use AlignScore to measure the alignment between the output summary and the input document.
-It serves as a quality indicator by evaluating the relevance and the overlap between the generated content with the source text.
-For translation tasks, we use COMET, as it captures both semantic adequacy and fluency, ensuring that translations are accurate and linguistically appropriate (Rei et al., 2020).
-Generation Setup.
-We discuss the generation parameters, the decoding strategy and sample selection procedure in depth in Appendix D.
-In short, we report evaluation results in two distinct setups: greedy decoding and stochastic sampling with focus on the most probable sequence among the generated outputs (best-sample).
-These two setups offer the highest-quality outputs and are the most reasonable generation approaches in practice.
 5.2 Results
 Main results.
-Tables 1 and 2 show the PRR scores under the best‐sample and greedy generation setup.
-We calculate a single representative PRR for each task – question answering, neural machine translation (NMT), and summarization (SUM) – by averaging the results across all relevant datasets (e.g., TriviaQA, MMLU, CoQA, GSM8k for QA).
+Table 1 shows the PRR scores under the greedy generation setup (See Appendix B for Best Sample and MBR Sample).
+We report aggregated PRR for each type of task – question answering, neural machine translation (NMT), and summarization (SUM) – by averaging the results across all relevant datasets (e.g., TriviaQA, MMLU, CoQA, GSM8k for QA).
 This aggregated score provides a concise measure of the performance for each model for each task.
-Detailed results for each dataset separately can be found in Appendix D.
-We can see that our CoCoA methods are the best across all tasks and models in our experiments.
-They outperform existing consistency‐based and hybrid state‐of‐the‐art approaches, like Semantic Entropy and SAR.
-In addition, our proposed CoCoA approach consistently surpasses the baseline UE metrics: for example, $\texttt{CoCoA}_{PPL}$ outperforms standard Perplexity, illustrating the advantage of combining token‐level confidence with semantic consistency.
-This pattern holds for other information‐based metrics as well, demonstrating that using the consistency between multiple sampled outputs reliably enhances uncertainty quantification.
+Detailed results for each dataset separately can be found in Appendix E.
+We can see that our CoCoA methods are the best across all tasks and models.
+They outperform existing consistency-based and hybrid state-of-the-art approaches, like Semantic Entropy and SAR.
+In addition, the proposed CoCoA approach consistently surpasses the baseline UQ measures: for example, $\texttt{CoCoA}_{PPL}$ outperforms standard Perplexity, illustrating the advantage of combining token-level confidence with semantic consistency.
+This pattern holds for other information-based metrics as well, demonstrating that using the consistency between multiple sampled outputs reliably enhances uncertainty quantification.
 Ablation study.
-As a part of our ablation study (see Appendix B.1), we evaluate the performance of the average dissimilarity component ($u_{*}^{\text{cons}}$) independently to assess its effectiveness as a standalone uncertainty measure and to investigate whether it could potentially outweigh the contribution of the information-based component ($u_{*}^{\text{info}}$) in the enriched uncertainty measure.
-This evaluation enables us to isolate and better understand the complementary roles and the relative importance of each component.
-Our experiments demonstrate that the combination of consistency- and confidence-based metrics outperforms the pure consistency-based measure on a vast majority of tasks.
-Notably, in the few cases where the pure consistency messure outperforms the combined approach, the performance difference is minimal.
-It is possible that a more suitable choice of a similarity measure or confidence-based metric for the task could further improve the performance.
-This leads us to the next part of our ablation study, where we investigate the impact of different similarity measures (see Appendix B.2).
+As a part of our ablation study, we investigate the impact of different similarity measures (see Appendix C.1).
 We find that for some tasks, the similarity score computed by the Cross-encoder does not yield optimal performance.
-For example, for question-answering tasks on CoQA and Trivia, NLI-derived similarity performs better than the Cross-encoder similarity and outperforms the pure consistency-based uncertainty discussed above.
-The next section of our ablation study focuses on alternative forms of combining model confidence $u_{*}^{\text{info}}$ and consistency $u_{*}^{\text{cons}}$ (see Apendix B.3).
-First, we consider an additive form of combining them: $\mathrm{U}_{\text{AdditiveCoCoA}}=u_{*}^{\text{info}}+u_{*}^{\text{cons}}$.
-The results show that this additive formulation does not perform as well compared to the multiplicative one.
+For example, for question-answering tasks on CoQA and TriviaQA, NLI-derived similarity performs better than the Cross-encoder similarity and outperforms the pure consistency-based uncertainty discussed above.
+The next section of our ablation study focuses on alternative forms of combining model confidence $u(\mathbf{y}_{*}\mid\mathbf{x})$ and consistency $\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})$; see Appendix C.2.
+First, we consider an additive form of combining them:
+$\widehat{U}_{\text{AdditiveCoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})+\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x}).$ (15)
+The results show that this additive formulation does not perform as well as the multiplicative one.
 The additive form tends to underemphasize the interaction between the two components, which is critical for capturing the nuanced relationships between confidence and consistency.
-We also consider an alternative formulation of the consistency term $u_{*}^{\text{cons}}$, as the average of the full pairwise dissimilarity.
-In this formulation, $u_{*}^{\text{cons}}$ represents the average inconsistency across all samples rather than focusing solely on the dissimilarity of the evaluated sequence with the other samples.
+We also consider an alternative formulation of the consistency term $\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})$, as the average of the full pairwise dissimilarity.
+In this formulation, $\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})$ represents the average inconsistency across all samples rather than focusing solely on the dissimilarity of the evaluated sequence with the other samples.
 Our experiments demonstrate that this formulation is not very strong.
 By distributing the consistency computation across all samples, it loses focus on the specific sequence being evaluated.
-Lastly, in Appendix B.3, we also consider alternative formulations of the information-based metric that do not rely on logarithmic transformations.
+Lastly, in Appendix C.2, we also consider alternative formulations of the information-based metric that do not rely on logarithmic transformations.
 While we primarily use logarithms due to their numerical stability, we explore an alternative approach by converting these values back to probabilities and analyzing their impact on uncertainty estimation.
@@
 Task and Domain Dependency.
-Our method relies on both an information-based confidence score and a semantic similarity function.
-The effectiveness of each can vary across models, tasks, and domains.
-For open-ended tasks with multiple equally valid outputs (e.g., creative generation), consistent rephrasing may inflate the perceived certainty.
-Conversely, in domains that demand highly precise factual or logical correctness (e.g., math problem solving), small deviations in reasoning can lead to large outcome differences that are not fully captured by a generic similarity measure.
-Adapting both the confidence measure and the similarity function to specific domains or prompt types is an important direction for future work.
+CoCoA relies on an information-based confidence score and a semantic similarity function, whose effectiveness can vary across models, tasks, and domains.
+In open-ended tasks like creative generation, producing diverse outputs is expected; on the other hand, tasks requiring precise reasoning can be sensitive to subtle errors that generic similarity metrics may miss.
+Adapting these components to specific domains remains an important direction for future work.
 Limited Sample Size.
-CoCoA estimates the model’s consistency by sampling multiple outputs and comparing them.
-In practice, generating a large number of samples can be computationally expensive and may increase the inference latency.
-Consequently, our experiments (like many sampling-based approaches) rely on relatively small sample sets.
-Although even a handful of samples can provide a meaningful estimate of consistency, it may not fully capture the diversity of the underlying distribution for certain tasks or for more complex prompts.
+CoCoA estimates consistency by sampling multiple outputs.
+Generating many samples is computationally expensive, so our experiments—like most sampling-based methods—use relatively small sets.
+While even a few samples can yield meaningful estimates, they may miss the full diversity of model outputs for complex prompts.
 Quality Metric.
 Finally, the CoCoA’s performance assessment depends on quality metrics (e.g., COMET for machine translation, and Accuracy for QA) that may not capture every nuance of textual outputs.
-Automatic metrics can have blind spots, particularly in evaluating coherence, factual correctness, or subtle aspects of style.
 Further refining or extending quality metrics to account for deeper reasoning, factual faithfulness, and stylistic appropriateness would better align uncertainty scores with real-world perceptions of model correctness.
@@
 We presented CoCoA, a unified approach that integrates Confidence and Consistency for uncertainty quantification in LLMs.
-By combining token-level confidence scores with semantic similarity between multiple sampled outputs, CoCoA offers a more holistic view of uncertainty than either approach alone.
+By combining confidence scores with semantic similarity between multiple sampled outputs, CoCoA offers a more holistic view of uncertainty than either approach alone.
 In extensive evaluations on question answering, summarization, and translation, our approach outperformed existing baselines and state-of-the-art UQ methods.
@@
 These include incorporating more adaptive sampling strategies that efficiently capture the model output space, refining semantic similarity functions for domain-specific tasks, and improving calibration techniques to strengthen the confidence metrics of the model.
-Impact Statement
-As large language models (LLMs) are increasingly integrated into everyday applications, ensuring the robustness of their outputs is essential, particularly in high-stakes domains such as healthcare and the legal sector.
-In addition to flagging low-quality responses, clear confidence estimates can help reduce overreliance on automated systems and foster safer, more responsible LLM usage.
-Here, we introduced CoCoA, a new framework that enhances uncertainty quantification in LLMs.
-While CoCoA improves reliability, its use requires careful domain adaptation to align with specific task requirements.
 Appendix A Decoding Strategy and Sample Selection
@@
 Table 2 shows average values of corresponding quality metrics for all combinations of models and datasets.
-We considered 4 approaches for the selection of output that subsequently is used to calculate the quality of generation.
-  • Greedy decoding produces single output by selecting top-1 candidate token at each generation step, thus not further selection of sample is needed.
+We considered 4 approaches for the selection of output that subsequently is used to calculate the quality of generation:
+  • Greedy decoding produces single output by selecting top-1 candidate token at each generation step, thus no further selection of sample is needed.
   • Random sample corresponds to the case where random output is selected among the number of samples produced by repeatedly prompting the model with the same question.
 In practice we use first generated sample, highlighting model performance when stochastic decoding is done only once.
-  • Best (normalized) sample selects the output with highest model-assigned (length-normalized) probability among several sampled outputs.
+  • Most Probable sample selects the output with highest model-assigned probability among several sampled outputs.
+  • MBR sample selects the output with highest average consistency with respect to other sampled outputs.
 We note that selecting a random sample from the model outputs incurs a significant drop in the quality of results on several datasets, most prominently on GSM8k.
-Based on these observations, we evaluate the efficacy of UE on two setups: greedy decoding and stochastic sampling with a focus on the highest-probability sample.
-In all experiments, we performed stochastic sampling with temperature $t=1.0$, top-k equal to 50, and top-p equal to $1.0$.
-Dataset Metric Llama Mistral Falcon
-Greedy Random Best Best Norm Greedy Random Best Best Norm Greedy Random Best Best Norm
-Trivia Algin Score 0.686 0.625 0.687 0.689 0.743 0.655 0.750 0.751 0.557 0.473 0.568 0.562
-MMLU Accuracy 0.590 0.430 0.597 0.587 0.633 0.558 0.632 0.632 0.713 0.639 0.713 0.713
-CoQa Algin Score 0.499 0.359 0.529 0.462 0.574 0.403 0.591 0.528 0.512 0.408 0.560 0.471
-GSM8k Accuracy 0.548 0.234 0.261 0.432 0.382 0.169 0.190 0.290 0.776 0.313 0.205 0.593
-Xsum Align Score 0.848 0.608 0.825 0.830 0.803 0.578 0.775 0.777 0.842 0.734 0.782 0.809
-WMT14FrEn Comet 0.863 0.819 0.852 0.859 0.863 0.812 0.830 0.855 0.867 0.833 0.857 0.863
-WMT19DeEn Comet 0.870 0.816 0.854 0.860 0.864 0.805 0.836 0.851 0.846 0.807 0.826 0.838
+Based on these observations, we evaluate the efficacy of UQ in three setups: greedy decoding, stochastic sampling with a focus on the most probable sample and MBR decoding.
+Regardless of the way the main model response was obtained, responses that were used to quantify consistency in (8) were generated via repeated stochastic sampling.
+In all the experiments, where stochastic sampling is involved, it was performed with temperature $T=1.0$, top-k equal to 50, and top-p equal to $1.0$.
+Sample Metric Greedy Random Sample Most Probable Sample MBR Sample
+Falcon7b-Base
+CoQA Align Score 0.793 0.706 0.783 0.793
+GSM8k Accuracy 0.776 0.313 0.205 0.419
+MMLU Accuracy 0.715 0.638 0.715 0.699
+TriviaQA Align Score 0.557 0.473 0.568 0.559
+WMT14FrEn Comet 0.867 0.833 0.857 0.855
+WMT19DeEn Comet 0.846 0.807 0.826 0.828
+XSUM Align Score 0.842 0.734 0.782 0.801
+Llama8b-Base
+CoQA Align Score 0.756 0.671 0.743 0.766
+GSM8k Accuracy 0.548 0.234 0.261 0.346
+MMLU Accuracy 0.570 0.368 0.577 0.469
+TriviaQA Align Score 0.686 0.625 0.687 0.699
+WMT14FrEn Comet 0.863 0.819 0.852 0.844
+WMT19DeEn Comet 0.870 0.816 0.854 0.845
+XSUM Align Score 0.848 0.608 0.825 0.703
+Mistral7b-Base
+CoQA Align Score 0.793 0.695 0.777 0.790
+GSM8k Accuracy 0.382 0.169 0.190 0.259
+MMLU Accuracy 0.632 0.552 0.632 0.612
+TriviaQA Align Score 0.743 0.655 0.750 0.739
+WMT14FrEn Comet 0.863 0.812 0.830 0.841
+WMT19DeEn Comet 0.864 0.805 0.836 0.837
+XSUM Align Score 0.803 0.578 0.775 0.665
 Table 2: Base quality metrics for models for different evaluated sequence choice.
-Appendix B Ablation
-B.1 Average Dissimilarity as Uncertainty Measure
-Tables 4 and 5 report PRRs of CoCoA-family methods with uncertainty estimates based solely on average dissimilarity of samples, as proposed in equation (8) We observe that it is still widely beneficial to synthesize consistency of outputs with model confidence, even when limiting consistency evaluation to the particular sample to be scored.
-Method Dataset
-XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
-Mistral7b-Base
-AveDissimilarity 0.051 0.285 0.500 0.379 0.647 0.423 0.435
-$\text{CoCoA}_{MSP}$ 0.330 0.396 0.598 0.383 0.670 0.466 0.517
-$\text{CoCoA}_{PPL}$ 0.286 0.375 0.568 0.369 0.674 0.466 0.467
-$\text{CoCoA}_{MTE}$ 0.288 0.374 0.564 0.355 0.673 0.447 0.491
-Llama8b-Base
-AveDissimilarity 0.024 0.389 0.453 0.375 0.614 0.392 0.368
-$\text{CoCoA}_{MSP}$ 0.378 0.456 0.582 0.349 0.597 0.485 0.372
-$\text{CoCoA}_{PPL}$ 0.387 0.448 0.514 0.338 0.593 0.452 0.433
-$\text{CoCoA}_{MTE}$ 0.380 0.446 0.511 0.337 0.601 0.402 0.447
-Falcon7b-Base
-AveDissimilarity 0.226 0.337 0.496 0.408 0.656 0.485 0.426
-$\text{CoCoA}_{MSP}$ 0.257 0.433 0.578 0.396 0.684 0.529 0.436
-$\text{CoCoA}_{PPL}$ 0.229 0.436 0.580 0.406 0.677 0.529 0.478
-$\text{CoCoA}_{MTE}$ 0.228 0.439 0.577 0.395 0.685 0.517 0.510
-Table 4: Comparison of PRRs of CoCoA-family methods with similarity of greedy output and other samples taken in isolation.
-Method Dataset
-XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
-Mistral7b-Base
-AveDissimilarity 0.071 0.670 0.708 0.405 0.614 0.423 0.846
-$\text{CoCoA}_{MSP}$ 0.366 0.712 0.730 0.430 0.644 0.466 0.562
-$\text{CoCoA}_{PPL}$ 0.372 0.735 0.755 0.402 0.648 0.466 0.937
-$\text{CoCoA}_{MTE}$ 0.373 0.708 0.732 0.373 0.645 0.447 0.935
-Llama8b-Base
-AveDissimilarity 0.030 0.473 0.598 0.395 0.600 0.353 0.795
-$\text{CoCoA}_{MSP}$ 0.484 0.529 0.685 0.384 0.587 0.452 0.513
-$\text{CoCoA}_{PPL}$ 0.450 0.544 0.689 0.364 0.573 0.422 0.925
-$\text{CoCoA}_{MTE}$ 0.451 0.520 0.638 0.346 0.582 0.363 0.900
-Falcon7b-Base
-AveDissimilarity 0.282 0.491 0.651 0.416 0.627 0.484 0.979
-$\text{CoCoA}_{MSP}$ 0.210 0.564 0.698 0.428 0.659 0.530 0.498
-$\text{CoCoA}_{PPL}$ 0.310 0.579 0.717 0.415 0.644 0.530 1.000
-$\text{CoCoA}_{MTE}$ 0.289 0.551 0.678 0.402 0.646 0.517 0.998
-Table 5: Comparison of PRRs of CoCoA-family methods with similarity of samples with best sample taken in isolation.
-B.2 Choice of Similarity Function
-For sample consistency estimation, one could come up with a variety of similarity functions $g(\mathbf{y},\mathbf{y}^{\prime})$.
+Appendix B Results Summary for Most Probable Sample and MBR Sample strategies
+Tables 3 and 4 report the PRR scores under the Most Probable Sample and MBR Sample generation setups, as discussed in Section  5.2.
+We observe that CoCoA-family methods, and their Supervised versions, consistently improve base uncertainty estimators.
+Metric Llama8b-Base Mistral7b-Base Falcon7b-Base
+QA NMT SUM QA NMT SUM QA NMT SUM
+MCSE 0.356 0.380 0.192 0.453 0.406 0.162 0.460 0.409 0.128
+MCNSE 0.380 0.429 0.186 0.466 0.489 0.196 0.530 0.424 0.153
+Semantic Entropy 0.396 0.411 0.194 0.482 0.438 0.164 0.479 0.440 0.134
+DegMat 0.422 0.342 0.191 0.465 0.425 0.205 0.543 0.386 0.177
+EigValLaplacian 0.388 0.274 0.190 0.426 0.366 0.197 0.498 0.336 0.174
+SAR 0.478 0.506 0.159 0.542 0.576 0.175 0.590 0.488 0.193
+Consistency 0.535 0.536 0.030 0.572 0.689 0.071 0.626 0.571 0.282
+MSP 0.395 0.376 0.464 0.444 0.252 0.330 0.343 0.381 0.099
+$\text{CoCoA}_{MSP}$ 0.484 $\uparrow$ 0.607 $\uparrow$ 0.484 $\uparrow$ 0.526 $\uparrow$ 0.721 $\uparrow$ 0.366 $\uparrow$ 0.529 $\uparrow$ 0.631 $\uparrow$ 0.210 $\uparrow$
+PPL 0.532 0.563 0.458 0.587 0.686 0.365 0.627 0.589 0.275
+$\text{CoCoA}_{PPL}$ 0.571 $\uparrow$ 0.617 $\uparrow$ 0.450 0.613 $\uparrow$ 0.745 $\uparrow$ 0.372 $\uparrow$ 0.647 $\uparrow$ 0.648 $\uparrow$ 0.310 $\uparrow$
+MTE 0.476 0.469 0.449 0.559 0.637 0.350 0.602 0.492 0.186
+$\text{CoCoA}_{MTE}$ 0.547 $\uparrow$ 0.579 $\uparrow$ 0.451 $\uparrow$ 0.600 $\uparrow$ 0.720 $\uparrow$ 0.373 $\uparrow$ 0.641 $\uparrow$ 0.614 $\uparrow$ 0.289 $\uparrow$
+Table 3: Results for Evaluated Sequence – Most Probable Sample: Mean PRR across datasets for each task.
+The best performing method is in bold, and the second-best is underscored.
+Arrows indicate improvement in CoCoA over the base version.
+Metric Llama8b-Base Mistral7b-Base Falcon7b-Base
+QA NMT SUM QA NMT SUM QA NMT SUM
+MCSE 0.402 0.297 0.158 0.474 0.330 0.237 0.437 0.346 0.076
+MCNSE 0.392 0.389 0.124 0.441 0.435 0.192 0.446 0.440 0.078
+Semantic Entropy 0.443 0.316 0.159 0.507 0.352 0.236 0.458 0.373 0.077
+DegMat 0.478 0.290 0.047 0.471 0.322 0.124 0.499 0.371 -0.018
+EigValLaplacian 0.434 0.224 0.037 0.437 0.279 0.115 0.460 0.326 -0.022
+SAR 0.509 0.456 0.187 0.532 0.448 0.266 0.531 0.478 0.083
+Consistency 0.504 0.365 0.235 0.518 0.367 0.284 0.538 0.410 0.099
+MSP 0.415 0.404 0.266 0.447 0.433 0.312 0.405 0.376 0.170
+$\text{CoCoA}_{MSP}$ 0.527 $\uparrow$ 0.496 $\uparrow$ 0.293 $\uparrow$ 0.553 $\uparrow$ 0.515 $\uparrow$ 0.348 $\uparrow$ 0.500 $\uparrow$ 0.526 $\uparrow$ 0.154
+PPL 0.408 0.421 0.164 0.432 0.452 0.273 0.445 0.438 0.100
+$\text{CoCoA}_{PPL}$ 0.532 $\uparrow$ 0.504 $\uparrow$ 0.215 $\uparrow$ 0.554 $\uparrow$ 0.514 $\uparrow$ 0.328 $\uparrow$ 0.552 $\uparrow$ 0.537 $\uparrow$ 0.109 $\uparrow$
+MTE 0.467 0.461 0.148 0.515 0.496 0.219 0.490 0.521 0.069
+$\text{CoCoA}_{MTE}$ 0.542 $\uparrow$ 0.479 $\uparrow$ 0.201 $\uparrow$ 0.574 $\uparrow$ 0.486 0.290 $\uparrow$ 0.562 $\uparrow$ 0.527 $\uparrow$ 0.081 $\uparrow$
+Table 4: Results for Evaluated Sequence - MBR Decoding: Mean PRR across datasets for each task.
+The best performing method is in bold, and the second-best is underscored.
+Arrows indicate improvement in CoCoA over the base version.
+Appendix C Ablation
+C.1 Choice of Similarity Function
+For sample consistency estimation, one could come up with a variety of similarity functions $s(\mathbf{y},\mathbf{y}^{\prime})$.
 We perform a comparison of the effectiveness of CoCoA-family methods using several such functions.
 We consider the following functions:
-  • AlignScore (Zha et al., 2023) with AlignScore-large model
-  • RougeL (Lin, 2004)
-  • NLI (He et al., 2021) based on microsoft/deberta-large-mnli model
-  • CrossEncoder (Liu et al., 2019) based on cross-encoder/stsb-roberta-large model.
-Tables 6 and 7 report these results.
+  • AlignScore Zha et al. [2023] with AlignScore-large model;
+  • RougeL Lin [2004];
+  • NLI He et al. [2021] based on microsoft/deberta-large-mnli model;
+  • CrossEncoder Liu et al. [2019] based on cross-encoder/stsb-roberta-large model.
+Tables 5, 6 and 7 report these results.
 There exists a considerable variation of relative effectiveness of proposed methods with various similarity function choices, depending on a task at hand.
@@
 Method Dataset
-XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
+XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
 Mistral7b-Base
@@
 CrossEncoder 0.223 0.438 0.575 0.395 0.685 0.517 0.505
-Table 6: Comparison of PRRs of CoCoA-family methods with different choices of similarity function with greedy sample taken in isolation.
+Table 5: Comparison of PRRs of CoCoA-family methods with different choices of the similarity function. Main model response obtained by greedy decoding.
 Method Dataset
-XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
+XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
 Mistral7b-Base
@@
 CrossEncoder 0.289 0.551 0.678 0.402 0.646 0.517 0.998
-Table 7: Comparison of PRRs of CoCoA-family methods with different choices of similarity function with best sample taken in isolation.
-B.3 Different Ways of Combining Confidence and Consistency
+Table 6: Comparison of PRRs of CoCoA-family methods with different choices of similarity function. Main model response obtained by selecting the most probable sample.
+Method Dataset
+XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
+Mistral7b-Base
+$\text{CoCoA}_{MSP}$
+AlignScore 0.287 0.326 0.460 0.312 0.575 0.515 0.747
+RougeL 0.259 0.420 0.562 0.304 0.502 0.515 0.750
+NLI 0.278 0.375 0.506 0.350 0.589 0.512 0.766
+CrossEncoder 0.295 0.441 0.589 0.336 0.597 0.514 0.766
+$\text{CoCoA}_{PPL}$
+AlignScore 0.283 0.333 0.482 0.330 0.596 0.515 0.738
+RougeL 0.219 0.427 0.561 0.300 0.517 0.515 0.744
+NLI 0.219 0.390 0.508 0.334 0.602 0.512 0.723
+CrossEncoder 0.258 0.451 0.577 0.322 0.612 0.514 0.768
+$\text{CoCoA}_{MTE}$
+AlignScore 0.269 0.314 0.447 0.357 0.663 0.497 0.746
+RougeL 0.181 0.446 0.571 0.327 0.622 0.497 0.786
+NLI 0.182 0.370 0.482 0.348 0.686 0.499 0.704
+CrossEncoder 0.236 0.419 0.553 0.337 0.680 0.494 0.783
+Llama8b-Base
+$\text{CoCoA}_{MSP}$
+AlignScore 0.221 0.309 0.462 0.261 0.467 0.622 0.717
+RougeL 0.195 0.401 0.542 0.250 0.361 0.631 0.721
+NLI 0.222 0.311 0.463 0.311 0.480 0.580 0.769
+CrossEncoder 0.225 0.403 0.590 0.304 0.487 0.586 0.731
+$\text{CoCoA}_{PPL}$
+AlignScore 0.225 0.325 0.484 0.304 0.476 0.603 0.747
+RougeL 0.169 0.416 0.537 0.274 0.370 0.612 0.758
+NLI 0.180 0.334 0.460 0.319 0.488 0.558 0.761
+CrossEncoder 0.198 0.428 0.580 0.320 0.496 0.562 0.751
+$\text{CoCoA}_{MTE}$
+AlignScore 0.209 0.293 0.460 0.320 0.563 0.556 0.749
+RougeL 0.148 0.436 0.559 0.314 0.497 0.587 0.785
+NLI 0.156 0.298 0.445 0.340 0.592 0.460 0.740
+CrossEncoder 0.189 0.398 0.561 0.345 0.587 0.472 0.765
+Falcon7b-Base
+$\text{CoCoA}_{MSP}$
+AlignScore 0.279 0.307 0.473 0.312 0.594 0.574 0.460
+RougeL 0.202 0.424 0.487 0.307 0.566 0.574 0.447
+NLI 0.231 0.362 0.508 0.355 0.618 0.577 0.531
+CrossEncoder 0.248 0.461 0.590 0.336 0.617 0.575 0.472
+$\text{CoCoA}_{PPL}$
+AlignScore 0.278 0.320 0.534 0.350 0.607 0.574 0.617
+RougeL 0.190 0.428 0.556 0.333 0.571 0.574 0.632
+NLI 0.216 0.376 0.549 0.368 0.623 0.578 0.661
+CrossEncoder 0.242 0.462 0.613 0.345 0.624 0.575 0.662
+$\text{CoCoA}_{MTE}$
+AlignScore 0.277 0.302 0.522 0.382 0.657 0.556 0.598
+RougeL 0.175 0.472 0.573 0.364 0.642 0.556 0.617
+NLI 0.196 0.359 0.530 0.400 0.671 0.562 0.639
+CrossEncoder 0.237 0.454 0.600 0.373 0.666 0.557 0.650
+Table 7: Comparison of PRRs of CoCoA-family methods with different choices of similarity function. Main model response obtained by MBR decoding.
+C.2 Different Ways of Combining Confidence and Consistency
 We justify the particular form of equation (11) by considering alternative ways to combine sample-focused confidence with consistency estimation.
-Results are presented in Tables 8 and 9.
+Results are presented in Tables 89 and 10.
 In particular, we investigate the performance of the additive approach (AdditiveCoCoA):
-$\mathrm{U}_{AdditiveCoCoA}=u_{*}^{\text{info}}+u_{*}^{\text{cons}},$ (11)
+$U_{\text{AdditiveCoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})+\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x}),$ (16)
 and the same multiplicative combination, replacing sample-focused dissimilarity from (8) with the average of the full pairwise dissimilarity matrix (24):
-$\mathrm{U}_{FullSampleCoCoA}=u_{*}^{\text{info}}\cdot U_{Deg}.$ (12)
+$U_{\text{FullSampleCoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})\cdot U_{\text{DegMat}}(\mathbf{x}).$ (17)
+In addition, we present results where we compute $u(\mathbf{y}_{*}\mid\mathbf{x})$ using probability-based scores instead of log-likelihood, see equation (6).
+We refer to this variant as $U_{\text{ProbCoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})$.
+While the results are generally similar, the log-likelihood-based formulation offers greater numerical stability.
+Therefore, we adopt it as the default in the main paper.
 It is evident that on average the multiplicative form proposed in equation (10) with both confidence and consistency terms focused on a single sample is the better performing variant.
 Method Dataset
-XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
+XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
 Mistral7b-Base
@@
 $\text{CoCoA}_{MTE}$ 0.228 0.439 0.577 0.395 0.685 0.517 0.510
-Table 8: Comparison of PRRs of CoCoA-family methods with alternative formulations with greedy sample taken in isolation.
+Table 8: Comparison of PRRs of CoCoA-family methods with alternative formulations. Main model response obtained via greedy decoding.
 Method Dataset
-XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
+XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
 Mistral7b-Base
@@
 $\text{CoCoA}_{MTE}$ 0.289 0.551 0.678 0.402 0.646 0.517 0.998
-Table 9: Comparison of PRRs of CoCoA-family methods with alternative formulations of with best sample taken in isolation.
-Appendix C Detailed Description of Uncertainty Quantification Methods
+Table 9: Comparison of PRRs of CoCoA-family methods with alternative formulations. Main model response obtained by selecting the most probable sample.
+Method Dataset
+XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
+Mistral7b-Base
+$\text{AdditiveCoCoA}_{MSP}$ 0.259 0.382 0.488 0.231 0.388 0.512 0.711
+$\text{FullSampleCoCoA}_{MSP}$ 0.287 0.447 0.591 0.318 0.570 0.516 0.748
+$\text{ProbCoCoA}_{MSP}$ 0.258 0.374 0.473 0.149 0.099 -0.265 0.710
+$\text{CoCoA}_{MTE}$ 0.295 0.441 0.589 0.336 0.597 0.514 0.766
+$\text{AdditiveCoCoA}_{PPL}$ 0.234 0.430 0.560 0.255 0.517 0.507 0.721
+$\text{FullSampleCoCoA}_{PPL}$ 0.251 0.458 0.571 0.305 0.587 0.516 0.745
+$\text{ProbCoCoA}_{PPL}$ 0.094 -0.123 -0.143 -0.202 -0.404 -0.423 0.027
+$\text{CoCoA}_{PPL}$ 0.258 0.451 0.577 0.322 0.612 0.514 0.768
+$\text{AdditiveCoCoA}_{MTE}$ 0.201 0.468 0.585 0.296 0.669 0.501 0.760
+$\text{FullSampleCoCoA}_{MTE}$ 0.232 0.426 0.544 0.318 0.673 0.496 0.772
+$\text{CoCoA}_{MTE}$ 0.236 0.419 0.553 0.337 0.680 0.494 0.783
+Llama8b-Base
+$\text{AdditiveCoCoA}_{MSP}$ 0.197 0.345 0.465 0.177 0.287 0.571 0.676
+$\text{FullSampleCoCoA}_{MSP}$ 0.214 0.409 0.589 0.289 0.449 0.589 0.724
+$\text{ProbCoCoA}_{MSP}$ 0.196 0.337 0.449 0.102 0.053 0.434 0.673
+$\text{CoCoA}_{MTE}$ 0.225 0.403 0.590 0.304 0.487 0.586 0.731
+$\text{AdditiveCoCoA}_{PPL}$ 0.182 0.406 0.500 0.229 0.401 0.540 0.715
+$\text{FullSampleCoCoA}_{PPL}$ 0.186 0.433 0.563 0.304 0.461 0.565 0.746
+$\text{ProbCoCoA}_{PPL}$ 0.101 -0.097 -0.152 -0.175 -0.367 0.099 0.053
+$\text{CoCoA}_{PPL}$ 0.198 0.428 0.580 0.320 0.496 0.562 0.751
+$\text{AdditiveCoCoA}_{MTE}$ 0.160 0.437 0.554 0.302 0.560 0.477 0.748
+$\text{FullSampleCoCoA}_{MTE}$ 0.174 0.400 0.543 0.334 0.568 0.475 0.762
+$\text{CoCoA}_{MTE}$ 0.189 0.398 0.561 0.345 0.587 0.472 0.765
+Falcon7b-Base
+$\text{AdditiveCoCoA}_{MSP}$ 0.188 0.356 0.401 0.228 0.488 0.578 0.368
+$\text{FullSampleCoCoA}_{MSP}$ 0.237 0.468 0.586 0.325 0.593 0.576 0.453
+$\text{ProbCoCoA}_{MSP}$ 0.184 0.339 0.378 0.144 0.350 -0.345 0.364
+$\text{CoCoA}_{MTE}$ 0.248 0.461 0.590 0.336 0.617 0.575 0.472
+$\text{AdditiveCoCoA}_{PPL}$ 0.206 0.404 0.565 0.271 0.537 0.572 0.625
+$\text{FullSampleCoCoA}_{PPL}$ 0.231 0.464 0.606 0.337 0.598 0.577 0.650
+$\text{ProbCoCoA}_{PPL}$ 0.014 -0.167 -0.265 -0.187 -0.237 -0.416 -0.206
+$\text{CoCoA}_{PPL}$ 0.242 0.462 0.613 0.345 0.624 0.575 0.662
+$\text{AdditiveCoCoA}_{MTE}$ 0.191 0.488 0.612 0.353 0.644 0.564 0.617
+$\text{FullSampleCoCoA}_{MTE}$ 0.219 0.452 0.591 0.368 0.654 0.556 0.629
+$\text{CoCoA}_{MTE}$ 0.237 0.454 0.600 0.373 0.666 0.557 0.650
+Table 10: Comparison of PRRs of CoCoA-family methods with alternative formulations. Main model response obtained by MBR decoding.
+Appendix D Detailed Description of Uncertainty Quantification Methods
 In this section, we provide a detailed description of the uncertainty quantification methods used in this study.
-C.1 Information-Based Methods
+D.1 Information-Based Methods
 Information-based methods are commonly used to estimate uncertainty by analyzing the probability distributions of tokens within a given output.
@@
 Thus, uncertainty is quantified by calculating the probability of the sequence with the highest likelihood, under the assumption that the model is most confident in this output.
+In its simplest form it is given by (6).
+An equivalent (in terms of ordering of predictions for different inputs) but more numerically stable formulation is logarithmic.
 It is defined as:
-$U_{MSP}(\mathbf{y}\mid\mathbf{x},\bm{\theta})=-\log P(\mathbf{y}\mid\mathbf{x}).$ (13)
-Perplexity (PPL) is another widely used metric for estimating uncertainty in language models (Fomicheva et al., 2020).
+$U_{\text{MSP}}(\mathbf{y}\mid\mathbf{x})=-\log p(\mathbf{y}\mid\mathbf{x}).$ (18)
+Perplexity (PPL) is another widely used metric for estimating uncertainty in language models [Fomicheva et al., 2020].
 It measures the model’s confidence by evaluating the average likelihood of generating the sequence tokens:
-$U_{\mathrm{PPL}}(\mathbf{y},\mathbf{x})=-\frac{1}{L}\log P(\mathbf{y}\mid\mathbf{x}).$ (14)
-Mean Token Entropy takes a broader view of uncertainty by considering the token-level predictions across the entire sequence (Fomicheva et al., 2020).
+$U_{\mathrm{PPL}}(\mathbf{y}\mid\mathbf{x})=-\frac{1}{L}\log p(\mathbf{y}\mid\mathbf{x}).$ (19)
+Mean Token Entropy takes a broader view of uncertainty by considering the token-level predictions across the entire sequence [Fomicheva et al., 2020].
 Instead of evaluating the model’s confidence in a single output or individual token predictions, Mean Token Entropy calculates the average entropy of the token probability distributions for each token in the sequence:
-$U_{\mathcal{H}_{T}}(\mathbf{y},\mathbf{x})=\frac{1}{L}\sum_{l=1}^{L}\mathcal{H}(y_{l}\mid\mathbf{y}_{<l},\mathbf{x}),$ (15)
+$U_{\mathcal{H}_{T}}(\mathbf{y}\mid\mathbf{x})=\frac{1}{L}\sum_{l=1}^{L}\mathcal{H}(y_{l}\mid\mathbf{y}_{<l},\mathbf{x}),$ (20)
 where $\mathcal{H}(y_{l}\mid\mathbf{y}_{<l},\mathbf{x})$ is an entropy of the token distribution $p(y_{l}\mid\mathbf{y}_{<l},\mathbf{x})$.
-The TokenSAR method, introduced in (Duan et al., 2024), generalizes length-normalized log probability by computing a weighted average of the negative log probabilities of generated tokens, where weights are based on token relevance to the overall text.
-Using a similarity function $g(\cdot,\cdot)$ and token relevance function $R_{T}(y_{k},\mathbf{y},\mathbf{x})=1-g(\mathbf{x}\cup\mathbf{y},\mathbf{x}\cup\mathbf{y}\setminus y_{k})$, the uncertainty estimate is calculated as:
-$U_{\mathrm{TokenSAR}}(\mathbf{y},\mathbf{x})=-\sum_{l=1}^{L}\tilde{\mathrm{R}}_{T}(y_{l},\mathbf{y},\mathbf{x})\log P(y_{l}\mid\mathbf{y}_{<l},\mathbf{x}),$ (16)
+The TokenSAR method, introduced in Duan et al.
+[2024], generalizes length-normalized log probability by computing a weighted average of the negative log probabilities of generated tokens, where weights are based on token relevance to the overall text.
+Using a similarity function $s(\cdot,\cdot)$ and token relevance function $R_{T}(y_{k},\mathbf{y},\mathbf{x})=1-s(\mathbf{x}\cup\mathbf{y},\mathbf{x}\cup\mathbf{y}\setminus y_{k})$, the uncertainty estimate is calculated as:
+$U_{\mathrm{TokenSAR}}(\mathbf{y}\mid\mathbf{x})=-\sum_{l=1}^{L}\tilde{\mathrm{R}}_{T}(y_{l},\mathbf{y},\mathbf{x})\log p(y_{l}\mid\mathbf{y}_{<l},\mathbf{x}),$ (21)
 where
-$\tilde{\mathrm{R}}_{T}(y_{k},\mathbf{y},\mathbf{x})=\frac{\mathrm{R}_{T}(y_{k},\mathbf{y},\mathbf{x})}{\sum\nolimits_{l=1}^{L}\mathrm{R}_{T}(y_{l},\mathbf{y},\mathbf{x})}.$ (17)
+$\tilde{\mathrm{R}}_{T}(y_{k},\mathbf{y}\mid\mathbf{x})=\frac{\mathrm{R}_{T}(y_{k},\mathbf{y},\mathbf{x})}{\sum\nolimits_{l=1}^{L}\mathrm{R}_{T}(y_{l},\mathbf{y},\mathbf{x})}.$ (22)
 This measure is central for computing SAR uncertainty measure.
-C.2 Consistency-Based Methods
+D.2 Consistency-Based Methods
 Consistency-based methods assess the uncertainty of a language model by evaluating the semantic consistency of its predictions across multiple outputs for the same prompt.
 The core idea is that semantically similar outputs indicate higher confidence, while diverse or conflicting outputs suggest greater uncertainty.
-Since language models can express the same meaning in different surface forms, these methods construct a semantic similarity matrix $G=(g_{ij})$, where each entry represents the degree of similarity between pairs of responses.
+Since language models can express the same meaning in different surface forms, these methods construct a semantic similarity matrix $S=(s_{ij})$, where each entry represents the degree of similarity between pairs of responses.
 By clustering responses into groups with equivalent meanings, these methods provide a semantic measure of the model’s consistency.
-Lin et al.
-(2024) offers two similarity measures to evaluate the similarity of sequences.
-The first is the Jaccard similarity, which treats sequences as sets of words and calculates the proportion of shared words to the total number of unique words in both sequences: $g(\mathbf{y},\mathbf{y}^{\prime})=|\mathbf{y}\cap\mathbf{y}^{\prime}|/|\mathbf{y}\cup\mathbf{y}^{\prime}|$.
+The work Lin et al.
+[2024] offers two similarity measures to evaluate the similarity of sequences.
+The first is the Jaccard similarity, which treats sequences as sets of words and calculates the proportion of shared words to the total number of unique words in both sequences: $s(\mathbf{y},\mathbf{y}^{\prime})=|\mathbf{y}\cap\mathbf{y}^{\prime}|/|\mathbf{y}\cup\mathbf{y}^{\prime}|$.
 Natural Language Inference (NLI) provides another method for computing similarity between sequences.
-We use the DeBERTa-large NLI model (He et al., 2021), following Kuhn et al.
-(2023).
+We use the DeBERTa-large NLI model He et al.
+[2021], following Kuhn et al.
+[2023].
 For each pair of sequences, an NLI model predicts two probabilities: ${p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})$, indicating entailment, and ${p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$, indicating contradiction.
-Similarity is then defined as either $g_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})={p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})$ or $g_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})=1-{p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$.
-Among the simplest consistency-based approaches are the Number of Semantic Sets and the Sum of Eigenvalues of the Graph Laplacian (Lin et al., 2024).
+Similarity is then defined as either $s_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})={p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})$ or $s_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})=1-{p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$.
+Among the simplest consistency-based approaches are the Number of Semantic Sets and the Sum of Eigenvalues of the Graph Laplacian [Lin et al., 2024].
 Number of Semantic Sets estimates how many distinct “meanings” the model produces by clustering its outputs with an NLI model.
 The number of semantic sets is initially equal to the total number of generated answers, $M$.
-Two sentences are grouped into the same cluster if the following conditions are satisfied: ${p}_{\text{entail}}(\mathbf{y}^{i},\mathbf{y}^{j})>{p}_{\text{contra}}(\mathbf{y}^{i},\mathbf{y}^{j})\quad\text{and}\quad{p}_{\text{entail}}(\mathbf{y}^{j},\mathbf{y}^{i})>{p}_{\text{contra}}(\mathbf{y}^{j},\mathbf{y}^{i})$.
+Two sentences are grouped into the same cluster if the following conditions are satisfied: ${p}_{\text{entail}}\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}>{p}_{\text{contra}}\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}\quad\text{and}\quad{p}_{\text{entail}}\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(i)}\bigr{)}>{p}_{\text{contra}}\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(i)}\bigr{)}$.
 This computation is performed for all pairs of answers, and the final number of distinct clusters is denoted by $U_{\text{NumSemSets}}$.
 Sum of Eigenvalues of the Graph Laplacian examines global diversity: it constructs a similarity matrix among the sampled outputs and computes a continuous uncertainty score from the eigenvalues of the Laplacian of that similarity graph.
-Lin et al.
-(2024) proposes computing an averaged similarity matrix as $g_{ij}=\bigl{(}g\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}+g\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(i)}\bigr{)}\bigr{)}/2$.
-The Laplacian for the matrix $G$ is defined as $L=I-D^{-\frac{1}{2}}GD^{-\frac{1}{2}}$, where $D$ is a diagonal matrix with elements $D_{ii}=\sum_{j=1}^{M}g_{ij}$.
+The work Lin et al.
+[2024] proposes computing an averaged similarity matrix as $s_{ij}=\bigl{(}s\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}+s\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(i)}\bigr{)}\bigr{)}/2$.
+The Laplacian for the matrix $S$ is defined as $L=I-D^{-\frac{1}{2}}SD^{-\frac{1}{2}}$, where $D$ is a (diagonal) degree matrix with elements $D_{ii}=\sum_{j=1}^{M}s_{ij}$.
 Consequently, the following formula is derived:
-$U_{\text{EigV}}=\sum_{i=1}^{M}\max(0,1-\lambda_{i}).$ (18)
+$U_{\text{EigV}}(\mathbf{x})=\sum_{i=1}^{M}\max(0,1-\lambda_{i}(\mathbf{x})).$ (23)
 Both Number of Semantic Sets and Sum of Eigenvalues of the Graph Laplacian effectively capture overall variation in generated text but cannot produce an individual uncertainty score for each output.
-To address this, Lin et al.
-(2024) proposes to use the diagonal Degree Matrix $D$ which represents the total similarity of each answer with all others.
-The corrected trace of $D$ provides an average pairwise distance between answers, and uncertainty is computed as:
-$U_{\mathrm{DegMat}}=1-\mathrm{trace}(D)/M^{2}.$ (19)
-C.3 Information-Based Methods with Repeated Sampling
-In this section we detail methods that integrate model confidence with consistency.
+To address this, the work Lin et al.
+[2024] proposes to use the diagonal Degree Matrix $D(\mathbf{x})$ which represents the total similarity of each answer with all others.
+The corrected trace of $D(\mathbf{x})$ provides an average pairwise distance between answers, and uncertainty is computed as:
+$U_{\mathrm{DegMat}}(\mathbf{x})=1-\mathrm{trace}(D(\mathbf{x}))/M^{2}.$ (24)
+D.3 Information-Based Methods with Repeated Sampling
+The natural idea is to somehow benefit from having multiple samples from the model while using important information contained in the output probabilities estimated by an LLM.
+Below, we examine several approaches that have sought to achieve this.
+Averaging uncertainties.
 We can compute the entropy on the sequence level $\mathbb{E}\bigl{[}-\log p(\mathbf{y}\mid\mathbf{x})\bigr{]}$, where the expectation is taken over the sequences $\mathbf{y}$ randomly generated from the distribution $p(\mathbf{y}\mid\mathbf{x})$.
@@
 In practice, we can use Monte-Carlo integration, i.e.
-generate several sequences $\mathbf{y}^{(i)},\,i=1,\dots,M$ via random sampling and compute Monte Carlo Sequence Entropy:
+generate several sequences $\mathbf{y}^{(i)},\,i=1,\dots,M$ via random sampling and compute Monte Carlo Sequence Entropy [Kuhn et al., 2023]:
 $U_{\mathcal{H}_{S}}(\mathbf{x})=-\frac{1}{M}\sum_{i=1}^{M}\log p(\mathbf{y}^{(i)}\mid\mathbf{x}).$ (25)
 We can replace $p(\mathbf{y}^{(i)}\mid\mathbf{x})$ with its length-normalized version $\bar{p}(\mathbf{y}^{(i)}\mid\mathbf{x})$ leading to a more reliable uncertainty measure in some cases.
-Semantic Entropy (Kuhn et al., 2023) addresses the issue of generated sequences with similar meanings but differing probabilities according to the model, which can heavily influence the resulting entropy value (20).
+While simple averaging represents a natural way to aggregate uncertainties, it has certain issues related to the nature of LLMs.
+First of all, in the vast majority of applications, an LLM-based system should produce a single output $\mathbf{y}_{*}$ for an input query.
+When we consider $U_{\mathcal{H}_{S}}(\mathbf{x})$ or other similar measure, we essentially perform averaging of uncertainties of different sequences, thus somewhat assessing the uncertainty related to the entire generative distribution $p(\mathbf{y}\mid\mathbf{x})$ for the input $\mathbf{x}$, but not for a particular generated sequence $\mathbf{y}_{*}$.
+This averaged uncertainty might not be adequate for this particular sequence and, remarkably, often performs worse than the uncertainty $u_{*}=U(\mathbf{y}_{*}\mid\mathbf{x})$, which is related solely to the output $\mathbf{y}_{*}$.
+Moreover, although intuitive, this naïve aggregation method assumes that all outputs contribute equally to the final uncertainty estimate, regardless of their semantic relationships.
+This can lead to inconsistencies when semantically equivalent outputs have varying uncertainty scores or when outputs with low similarity are treated as equally important.
+Semantically weighted averaging.
+Semantic Entropy Kuhn et al.
+[2023] addresses the issue of generated sequences with similar meanings but differing probabilities according to the model, which can heavily influence the resulting entropy value (25).
 The method clusters generated sequences $\mathbf{y}^{(i)},\,i=1,\dots,M$ into semantically homogeneous groups $\mathcal{C}_{k},~{}k=1,\dots,K$ (where $K\leq M$) using a bi-directional entailment algorithm.
@@
 where $\hat{p}_{k}(\mathbf{x})=\sum_{\mathbf{y}\in\mathcal{C}_{k}}p(\mathbf{y}\mid\mathbf{x})$ represents the aggregated probability for cluster $\mathcal{C}_{k}$.
-SentenceSAR (Duan et al., 2024) enhances the probability of sentences that are more relevant.
-It uses a sentence relevance measure $g\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(k)}\bigr{)}$ to evaluate the relevance of $\mathbf{y}^{(j)}$ with respect to $\mathbf{y}^{(k)}$.
+SentenceSAR Duan et al.
+[2024] enhances the probability of sentences that are more relevant.
+It uses a sentence relevance measure $s\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(k)}\bigr{)}$ to evaluate the relevance of $\mathbf{y}^{(j)}$ with respect to $\mathbf{y}^{(k)}$.
 SentenceSAR is calculated as:
@@
 where $t$ is a temperature parameter used to control the scale of shifting to relevance, and
-$\mathrm{R}_{S}(\mathbf{y}^{(j)},\mathbf{x})\!=\sum_{k\neq j}g\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(k)}\bigr{)}P\bigl{(}\mathbf{y}^{(k)}\mid\mathbf{x}\bigr{)}.$ (23)
-The combination of SentenceSAR and TokenSAR results in a unified method called SAR (Duan et al., 2024).
+$\mathrm{R}_{S}(\mathbf{y}^{(j)},\mathbf{x})\!=\sum_{k\neq j}s\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(k)}\bigr{)}p\bigl{(}\mathbf{y}^{(k)}\mid\mathbf{x}\bigr{)}.$ (28)
+The combination of SentenceSAR and TokenSAR results in a unified method called SAR Duan et al.
+[2024].
 In this approach, the generative probability $p(\mathbf{y}\mid\mathbf{x})$ in the SentenceSAR formula is replaced with the token-shifted probability $p^{\prime}(\mathbf{y}\mid\mathbf{x})=\exp\bigl{\{}-\mathrm{TokenSAR}(\mathbf{y},\mathbf{x})\bigr{\}}$, creating a comprehensive measure that integrates both sentence- and token-level adjustments.
-Appendix D Detailed Experimental Results
-In this section, we present detailed experimental results, which were used for computing values in Tables 1 and 2.
+The aggregation approaches like Semantic Entropy Kuhn et al. [2023] or SAR Duan et al. [2024] can be unified into a semantically-aware Generalized Monte Carlo uncertainty estimate, defined as
+$U_{\text{GMCU}}(\mathbf{x})=\frac{1}{M}\sum_{i=1}^{M}h\Biggl{(}\sum_{j=1}^{M}s\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}p\bigl{(}\mathbf{y}^{(j)}\mid\mathbf{x}\bigr{)}\Biggr{)}.$ (29)
+Here, the inner summation aggregates sequence probabilities $p\bigl{(}\mathbf{y}^{(j)}\mid\mathbf{x}\bigr{)}$ weighted by their semantic similarity to the $i$-th output, and the outer summation averages these contributions across all samples.
+The function $h(\cdot)$ provides an additional layer of flexibility, transforming the reweighted uncertainty scores, making the method a generalized framework for uncertainty quantification.
+Unfortunately, methods that fall under GMCU, while offering benefits, also inherit the aforementioned issues from both categories of methods.
+In particular, the outer summation in (29), similarly to the case of simple Monte Carlo averaging, often fails to outperform the uncertainty $U_{*}(\mathbf{x})=U(\mathbf{y}_{*}\mid\mathbf{x})$ of a single generated sequence $\mathbf{y}_{*}$.
+Appendix E Detailed Experimental Results
+Tables 3 and 1 presented PRR scores averaged over datasets corresponding to each task. Here we present expanded results for each dataset, see Tables 11, 12 and 13.
 Method Dataset
-XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
+XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
 Mistral7b-Base
@@
 Semantic Entropy 0.008 0.271 0.382 0.271 0.562 0.387 0.472
-SAR 0.094 0.353 0.517 0.313 0.644 0.419 0.471
 DegMat 0.137 0.229 0.382 0.336 0.646 0.410 0.299
 EigValLaplacian 0.132 0.207 0.328 0.301 0.624 0.398 0.241
+SAR 0.094 0.353 0.517 0.313 0.644 0.419 0.471
+Consistency Light -0.006 0.317 0.457 0.306 0.597 0.432 0.441
+Consistency 0.051 0.285 0.500 0.379 0.647 0.423 0.435
 MSP 0.287 0.315 0.451 0.326 0.628 0.474 0.471
 $\text{CoCoA}_{MSP}$ 0.330 0.396 0.598 0.383 0.670 0.466 0.517
+$\text{CoCoA}_{MSP}$ Light 0.309 0.380 0.580 0.351 0.674 0.480 0.506
 PPL 0.204 0.365 0.489 0.281 0.632 0.474 0.311
 $\text{CoCoA}_{PPL}$ 0.286 0.375 0.568 0.369 0.674 0.466 0.467
+$\text{CoCoA}_{PPL}$ Light 0.260 0.404 0.557 0.326 0.684 0.480 0.428
 MTE 0.182 0.392 0.484 0.243 0.619 0.456 0.350
 $\text{CoCoA}_{MTE}$ 0.288 0.374 0.564 0.355 0.673 0.447 0.491
+$\text{CoCoA}_{MTE}$ Light 0.254 0.415 0.550 0.303 0.672 0.465 0.458
 Llama8b-Base
@@
 Semantic Entropy 0.033 0.297 0.389 0.272 0.549 0.229 0.375
-SAR 0.077 0.427 0.483 0.311 0.595 0.352 0.398
 DegMat 0.081 0.250 0.355 0.353 0.622 0.342 0.309
 EigValLaplacian 0.079 0.198 0.278 0.332 0.604 0.292 0.273
+SAR 0.077 0.427 0.483 0.311 0.595 0.352 0.398
+Consistency Light -0.022 0.475 0.441 0.353 0.552 0.279 0.378
+Consistency 0.024 0.389 0.453 0.375 0.614 0.391 0.368
 MSP 0.328 0.342 0.456 0.277 0.526 0.508 0.324
 $\text{CoCoA}_{MSP}$ 0.378 0.456 0.582 0.349 0.597 0.485 0.372
+$\text{CoCoA}_{MSP}$ Light 0.358 0.448 0.556 0.340 0.598 0.504 0.353
 PPL 0.369 0.351 0.422 0.253 0.507 0.461 0.303
 $\text{CoCoA}_{PPL}$ 0.387 0.448 0.514 0.338 0.593 0.452 0.433
-MTE 0.357 0.357 0.408 0.239 0.497 0.350 0.326
-$\text{CoCoA}_{MTE}$ 0.380 0.446 0.511 0.337 0.601 0.402 0.447
+$\text{CoCoA}_{PPL}$ Light 0.382 0.483 0.491 0.319 0.597 0.467 0.398
+MTE 0.357 0.357 0.408 0.239 0.497 0.349 0.326
+$\text{CoCoA}_{MTE}$ 0.380 0.446 0.511 0.337 0.601 0.401 0.447
+$\text{CoCoA}_{MTE}$ Light 0.372 0.501 0.487 0.316 0.599 0.387 0.412
 Falcon7b-Base
@@
 Semantic Entropy 0.164 0.307 0.389 0.294 0.581 0.463 0.418
-SAR 0.144 0.398 0.517 0.381 0.649 0.508 0.387
 DegMat 0.201 0.274 0.431 0.407 0.651 0.480 0.395
 EigValLaplacian 0.201 0.229 0.394 0.381 0.645 0.454 0.358
+SAR 0.144 0.398 0.517 0.381 0.649 0.508 0.387
+Consistency Light 0.232 0.483 0.468 0.327 0.657 0.362 0.363
+Consistency 0.226 0.337 0.496 0.408 0.656 0.485 0.426
 MSP 0.201 0.312 0.400 0.321 0.662 0.539 0.377
 $\text{CoCoA}_{MSP}$ 0.257 0.433 0.578 0.396 0.684 0.529 0.436
+$\text{CoCoA}_{MSP}$ Light 0.242 0.463 0.566 0.349 0.691 0.522 0.405
 PPL 0.155 0.375 0.525 0.316 0.644 0.539 0.326
 $\text{CoCoA}_{PPL}$ 0.229 0.436 0.580 0.406 0.677 0.529 0.478
+$\text{CoCoA}_{PPL}$ Light 0.234 0.497 0.558 0.364 0.694 0.522 0.468
 MTE 0.152 0.409 0.537 0.291 0.633 0.533 0.367
 $\text{CoCoA}_{MTE}$ 0.228 0.439 0.577 0.395 0.685 0.517 0.510
-Table 10: Detailed experimental results with greedy sample taken in isolation.
+$\text{CoCoA}_{MTE}$ Light 0.234 0.518 0.549 0.345 0.693 0.491 0.500
+Table 11: Detailed experimental results. Main model response obtained by greedy decoding.
 Method Dataset
-XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
+XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
 Mistral7b-Base
@@
 Semantic Entropy 0.164 0.434 0.442 0.312 0.554 0.389 0.675
-SAR 0.175 0.563 0.590 0.347 0.620 0.421 0.780
 DegMat 0.205 0.439 0.410 0.376 0.618 0.410 0.454
 EigValLaplacian 0.197 0.388 0.344 0.342 0.600 0.399 0.361
+SAR 0.175 0.563 0.590 0.347 0.620 0.421 0.780
+Consistency 0.071 0.670 0.708 0.405 0.614 0.423 0.846
 MSP 0.330 0.212 0.291 0.388 0.607 0.476 0.307
@@
 Llama8b-Base
-MCSE 0.192 0.366 0.395 0.259 0.465 0.158 0.546
-MCNSE 0.186 0.377 0.480 0.239 0.484 0.165 0.634
-Semantic Entropy 0.194 0.371 0.451 0.286 0.528 0.213 0.559
-SAR 0.159 0.441 0.571 0.327 0.578 0.340 0.668
-DegMat 0.191 0.274 0.409 0.367 0.606 0.320 0.397
-EigValLaplacian 0.190 0.216 0.333 0.340 0.587 0.274 0.351
-MSP 0.464 0.339 0.413 0.304 0.514 0.483 0.281
+MCSE 0.192 0.366 0.395 0.258 0.465 0.158 0.545
+MCNSE 0.186 0.377 0.480 0.239 0.484 0.165 0.631
+Semantic Entropy 0.194 0.371 0.451 0.286 0.528 0.213 0.557
+DegMat 0.191 0.274 0.409 0.366 0.606 0.320 0.396
+EigValLaplacian 0.190 0.216 0.333 0.339 0.587 0.274 0.351
+SAR 0.159 0.441 0.571 0.327 0.578 0.340 0.667
+Consistency 0.030 0.473 0.598 0.394 0.600 0.353 0.793
+MSP 0.464 0.339 0.413 0.304 0.514 0.483 0.280
 $\text{CoCoA}_{MSP}$ 0.484 0.529 0.685 0.384 0.587 0.452 0.513
-PPL 0.458 0.504 0.622 0.294 0.483 0.441 0.912
-$\text{CoCoA}_{PPL}$ 0.450 0.544 0.689 0.364 0.573 0.422 0.925
-MTE 0.449 0.437 0.501 0.239 0.458 0.326 0.884
-$\text{CoCoA}_{MTE}$ 0.451 0.520 0.638 0.346 0.582 0.363 0.900
+PPL 0.458 0.504 0.622 0.293 0.483 0.441 0.911
+$\text{CoCoA}_{PPL}$ 0.450 0.544 0.689 0.364 0.573 0.422 0.924
+MTE 0.449 0.437 0.501 0.238 0.458 0.326 0.883
+$\text{CoCoA}_{MTE}$ 0.451 0.520 0.638 0.345 0.582 0.363 0.899
 Falcon7b-Base
@@
 Semantic Entropy 0.134 0.420 0.460 0.319 0.566 0.463 0.567
-SAR 0.193 0.455 0.521 0.385 0.642 0.509 0.826
 DegMat 0.177 0.350 0.422 0.422 0.637 0.480 0.633
 EigValLaplacian 0.174 0.289 0.382 0.393 0.622 0.454 0.522
+SAR 0.193 0.455 0.521 0.385 0.642 0.509 0.826
+Consistency 0.282 0.491 0.651 0.416 0.627 0.484 0.979
 MSP 0.099 0.385 0.378 0.369 0.638 0.540 -0.175
@@
 $\text{CoCoA}_{MTE}$ 0.289 0.551 0.678 0.402 0.646 0.517 0.998
-Table 11: Detailed experimental results with best sample taken in isolation.
+Table 12: Detailed experimental results. Main model response obtained by selecting most probable candidate among stochastically sampled responses.
+Method Dataset
+XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
+Mistral7b-Base
+MCSE 0.123 0.297 0.363 0.245 0.515 0.385 0.751
+MCNSE 0.105 0.385 0.485 0.256 0.552 0.406 0.548
+Semantic Entropy 0.127 0.310 0.395 0.272 0.580 0.439 0.735
+DegMat 0.244 0.280 0.365 0.358 0.673 0.461 0.392
+EigValLaplacian 0.243 0.238 0.319 0.334 0.654 0.443 0.318
+SAR 0.204 0.392 0.503 0.334 0.658 0.478 0.656
+Consistency 0.211 0.297 0.437 0.355 0.659 0.458 0.603
+MSP 0.259 0.381 0.486 0.216 0.346 0.514 0.711
+$\text{CoCoA}_{MTE}$ 0.295 0.441 0.589 0.336 0.597 0.514 0.766
+PPL 0.205 0.391 0.513 0.186 0.358 0.514 0.672
+$\text{CoCoA}_{PPL}$ 0.258 0.451 0.577 0.322 0.612 0.514 0.768
+MTE 0.156 0.443 0.550 0.234 0.598 0.508 0.717
+$\text{CoCoA}_{MTE}$ 0.236 0.419 0.553 0.337 0.680 0.494 0.783
+Llama8b-Base
+MCSE 0.089 0.240 0.354 0.220 0.444 0.297 0.647
+MCNSE 0.055 0.334 0.443 0.247 0.470 0.306 0.545
+Semantic Entropy 0.089 0.250 0.382 0.245 0.514 0.361 0.650
+DegMat 0.217 0.218 0.361 0.368 0.611 0.391 0.539
+EigValLaplacian 0.217 0.167 0.280 0.347 0.592 0.305 0.492
+SAR 0.125 0.393 0.520 0.345 0.578 0.472 0.642
+Consistency 0.146 0.290 0.440 0.387 0.603 0.428 0.599
+MSP 0.197 0.344 0.463 0.161 0.247 0.577 0.675
+$\text{CoCoA}_{MTE}$ 0.225 0.403 0.590 0.304 0.487 0.586 0.731
+PPL 0.168 0.382 0.460 0.161 0.253 0.549 0.668
+$\text{CoCoA}_{PPL}$ 0.198 0.428 0.580 0.320 0.496 0.562 0.751
+MTE 0.130 0.412 0.509 0.245 0.466 0.459 0.699
+$\text{CoCoA}_{MTE}$ 0.189 0.398 0.561 0.345 0.587 0.472 0.765
+Falcon7b-Base
+MCSE 0.188 0.331 0.362 0.274 0.560 0.455 0.460
+MCNSE 0.126 0.387 0.494 0.306 0.605 0.479 0.396
+Semantic Entropy 0.192 0.332 0.414 0.304 0.594 0.497 0.438
+DegMat 0.246 0.271 0.470 0.392 0.672 0.516 0.415
+EigValLaplacian 0.238 0.220 0.432 0.365 0.660 0.476 0.339
+SAR 0.184 0.412 0.543 0.370 0.664 0.553 0.536
+Consistency 0.235 0.323 0.498 0.376 0.665 0.523 0.588
+MSP 0.187 0.355 0.398 0.213 0.458 0.583 0.367
+$\text{CoCoA}_{MSP}$ 0.248 0.461 0.590 0.336 0.617 0.575 0.472
+PPL 0.177 0.371 0.504 0.201 0.439 0.583 0.556
+$\text{CoCoA}_{PPL}$ 0.242 0.462 0.613 0.345 0.624 0.575 0.662
+MTE 0.159 0.462 0.580 0.290 0.588 0.565 0.517
+$\text{CoCoA}_{MTE}$ 0.237 0.454 0.600 0.373 0.666 0.557 0.650
+Table 13: Detailed experimental results. Main model response obtained by MBR decoding.
+Appendix F Alternative Performance Metrics for PRR
+The choice of PRR as a UQ quality metric of choice is dictated by its ability to handle both continuous performance metrics, like Comet and AlignScore without the need for selecting arbitrary thresholds, as well as relative robustness to class imbalance in case of binary performance metric.
+However, PRR scores are calculated for a particular choice of underlying performance metric.
+For a truly comprehensive evaluation, we perform the same evaluation as in our main experimental run, but with alternative choice of performance metrics.
+Tables 14 and 15 report PRRs for these metrics on all models for NMT and QA tasks.
+MetricX Juraska et al.
+[2024] was used for NMT and GPT-as-a-judge (gpt-4o-2024-08-06) was used to score QA datasets.
+The following is the prompt used to facilitate GPT QA scoring:
+You are a text evaluator. The model was asked the following question:
+{question}
+The ’Generated’ text is a model’s response. The ’Target’ is the correct answer.
+If the generated answer correctly answers the question based on the target, return 1.
+If it is wrong, return 0.
+Respond ONLY with a single digit: 1 or 0.
+Generated: {model output}
+Target: {target sequence}
+NMT (MetricX) QA (GPT-4o)
+Method WMT14FrEn WMT19DeEn CoQA/Gpt TriviaQA/Gpt MMLU/Gpt GSM8k/Gpt
+Mistral7b-Base
+MCSE 0.127 0.212 0.226 0.478 0.338 0.453
+MCNSE 0.267 0.417 0.246 0.535 0.357 0.388
+Semantic Entropy 0.158 0.261 0.254 0.554 0.387 0.454
+CEDegMat 0.270 0.402 0.336 0.659 0.402 0.356
+SAR 0.302 0.451 0.333 0.660 0.419 0.456
+Semantic Density 0.291 0.366 0.056 0.688 0.335 0.195
+MSP 0.173 0.287 0.287 0.640 0.473 0.485
+$\text{CoCoA}_{MSP}$ 0.314 0.462 0.367 0.681 0.465 0.537
+PPL 0.316 0.448 0.220 0.645 0.473 0.345
+$\text{CoCoA}_{PPL}$ 0.361 0.512 0.331 0.687 0.465 0.499
+MTE 0.350 0.455 0.190 0.631 0.456 0.387
+$\text{CoCoA}_{MTE}$ 0.366 0.512 0.323 0.684 0.447 0.524
+Llama8b-Base
+MCSE 0.145 0.181 0.200 0.486 0.158 0.332
+MCNSE 0.319 0.361 0.183 0.530 0.170 0.324
+Semantic Entropy 0.157 0.235 0.239 0.552 0.203 0.361
+CEDegMat 0.287 0.367 0.350 0.621 0.299 0.300
+SAR 0.354 0.420 0.302 0.610 0.319 0.382
+Semantic Density 0.252 0.328 0.069 0.650 0.335 0.299
+MSP 0.176 0.274 0.228 0.564 0.469 0.313
+$\text{CoCoA}_{MSP}$ 0.329 0.451 0.332 0.625 0.449 0.365
+PPL 0.327 0.388 0.217 0.549 0.462 0.309
+$\text{CoCoA}_{PPL}$ 0.395 0.469 0.329 0.623 0.443 0.445
+MTE 0.353 0.388 0.214 0.543 0.358 0.329
+$\text{CoCoA}_{MTE}$ 0.394 0.472 0.333 0.634 0.393 0.455
+Falcon7b-Base
+MCSE 0.168 0.218 0.228 0.587 0.420 0.443
+MCNSE 0.304 0.399 0.265 0.640 0.442 0.323
+Semantic Entropy 0.201 0.276 0.262 0.623 0.463 0.448
+CEDegMat 0.287 0.428 0.380 0.712 0.476 0.347
+SAR 0.339 0.457 0.367 0.715 0.508 0.410
+Semantic Density 0.323 0.422 0.076 0.746 0.396 0.349
+MSP 0.176 0.275 0.279 0.737 0.539 0.391
+$\text{CoCoA}_{MSP}$ 0.335 0.483 0.364 0.763 0.529 0.459
+PPL 0.322 0.454 0.245 0.722 0.539 0.341
+$\text{CoCoA}_{PPL}$ 0.394 0.531 0.345 0.754 0.529 0.513
+MTE 0.370 0.455 0.228 0.707 0.533 0.385
+$\text{CoCoA}_{MTE}$ 0.403 0.529 0.341 0.762 0.516 0.551
+Table 14: PRRs for all models on QA and NMT tasks with alternative choice of performance metrics. Main model response obtained by greedy decoding.
+NMT (MetricX) QA (GPT-4o)
+Method WMT14FrEn WMT19DeEn CoQA/Gpt TriviaQA/Gpt MMLU/Gpt GSM8k/Gpt
+Mistral7b-Base
+MCSE 0.305 0.324 0.244 0.464 0.339 0.392
+MCNSE 0.360 0.466 0.252 0.516 0.358 0.188
+Semantic Entropy 0.336 0.366 0.272 0.536 0.389 0.420
+CEDegMat 0.468 0.485 0.339 0.623 0.404 0.239
+SAR 0.455 0.527 0.336 0.628 0.421 0.298
+Semantic Density 0.565 0.554 0.043 0.646 0.306 0.421
+MSP 0.169 0.252 0.311 0.608 0.474 0.460
+$\text{CoCoA}_{MSP}$ 0.607 0.623 0.376 0.647 0.466 0.550
+PPL 0.510 0.552 0.265 0.619 0.474 0.096
+$\text{CoCoA}_{PPL}$ 0.620 0.667 0.346 0.655 0.466 0.091
+MTE 0.487 0.499 0.211 0.604 0.457 0.167
+$\text{CoCoA}_{MTE}$ 0.581 0.638 0.324 0.653 0.447 0.151
+Llama8b-Base
+MCSE 0.284 0.271 0.191 0.451 0.155 0.302
+MCNSE 0.354 0.419 0.197 0.493 0.171 0.251
+Semantic Entropy 0.299 0.333 0.230 0.515 0.201 0.341
+CEDegMat 0.343 0.451 0.369 0.598 0.301 0.241
+SAR 0.412 0.497 0.315 0.584 0.323 0.311
+Semantic Density 0.403 0.452 0.065 0.638 0.286 0.396
+MSP 0.247 0.289 0.225 0.523 0.467 0.332
+$\text{CoCoA}_{MSP}$ 0.473 0.568 0.346 0.597 0.440 0.418
+PPL 0.437 0.517 0.238 0.498 0.459 0.185
+$\text{CoCoA}_{PPL}$ 0.510 0.614 0.335 0.588 0.433 0.199
+MTE 0.387 0.439 0.201 0.485 0.345 0.226
+$\text{CoCoA}_{MTE}$ 0.495 0.589 0.330 0.604 0.372 0.237
+Falcon7b-Base
+MCSE 0.269 0.341 0.261 0.553 0.421 0.159
+MCNSE 0.325 0.403 0.312 0.620 0.443 0.064
+Semantic Entropy 0.308 0.397 0.296 0.588 0.463 0.172
+CEDegMat 0.381 0.478 0.398 0.686 0.477 0.083
+SAR 0.389 0.483 0.398 0.685 0.509 0.152
+Semantic Density 0.434 0.518 0.047 0.661 0.377 0.299
+MSP 0.269 0.311 0.340 0.683 0.540 0.104
+$\text{CoCoA}_{MSP}$ 0.465 0.621 0.407 0.705 0.530 0.300
+PPL 0.430 0.553 0.308 0.663 0.540 0.156
+$\text{CoCoA}_{PPL}$ 0.526 0.655 0.376 0.691 0.530 0.128
+MTE 0.418 0.452 0.275 0.615 0.535 0.054
+$\text{CoCoA}_{MTE}$ 0.512 0.620 0.370 0.692 0.517 0.050
+Table 15: PRRs for all models on QA and NMT tasks with alternative choice of performance metrics.
+Main model response obtained by selecting the most probable candidate among stochastically sampled responses.
+Appendix G AUROC
+We strongly believe that due to the considerations presented in Appendix F, PRR is a superior metric to AUROC when comparing relative performance of UQ methods.
+However, we acknowledge that AUROC is widely used in modern UQ literature, so we opt to include AUROC results on QA datasets.
+The results are presented in Tables 16 and 17.
+All results here were obtained with GPT-as-a-judge correctness scoring.
+CoQA/Gpt TriviaQA/Gpt MMLU/Gpt GSM8k/Gpt
+Mistral7b-Base
+MCSE 0.642 0.784 0.737 0.716
+MCNSE 0.653 0.806 0.744 0.684
+Semantic Entropy 0.653 0.817 0.756 0.716
+CEDegMat 0.690 0.859 0.759 0.682
+SAR 0.694 0.862 0.770 0.717
+Semantic Density 0.531 0.871 0.713 0.576
+MSP 0.674 0.849 0.794 0.684
+$\text{CoCoA}_{MSP}$ 0.708 0.867 0.790 0.729
+PPL 0.636 0.855 0.794 0.628
+$\text{CoCoA}_{PPL}$ 0.690 0.871 0.790 0.714
+MTE 0.620 0.849 0.787 0.640
+$\text{CoCoA}_{MTE}$ 0.687 0.870 0.783 0.727
+Llama8b-Base
+MCSE 0.635 0.779 0.620 0.705
+MCNSE 0.630 0.799 0.629 0.689
+Semantic Entropy 0.655 0.815 0.652 0.713
+CEDegMat 0.708 0.845 0.703 0.691
+SAR 0.694 0.842 0.715 0.724
+Semantic Density 0.540 0.855 0.680 0.655
+MSP 0.647 0.816 0.787 0.669
+$\text{CoCoA}_{MSP}$ 0.700 0.847 0.776 0.718
+PPL 0.640 0.815 0.778 0.661
+$\text{CoCoA}_{PPL}$ 0.699 0.848 0.770 0.745
+MTE 0.633 0.817 0.726 0.667
+$\text{CoCoA}_{MTE}$ 0.702 0.854 0.746 0.753
+Falcon7b-Base
+MCSE 0.648 0.807 0.770 0.766
+MCNSE 0.660 0.831 0.779 0.709
+Semantic Entropy 0.669 0.830 0.788 0.761
+CEDegMat 0.718 0.870 0.784 0.722
+SAR 0.713 0.870 0.805 0.757
+Semantic Density 0.544 0.881 0.746 0.716
+MSP 0.681 0.866 0.820 0.731
+$\text{CoCoA}_{MSP}$ 0.719 0.884 0.815 0.774
+PPL 0.645 0.860 0.820 0.700
+$\text{CoCoA}_{PPL}$ 0.705 0.882 0.815 0.800
+MTE 0.630 0.854 0.818 0.721
+$\text{CoCoA}_{MTE}$ 0.704 0.886 0.810 0.814
+Table 16: AUROC for all models on QA tasks. Main model response was obtained by greedy decoding.
+CoQA/Gpt TriviaQA/Gpt MMLU/Gpt GSM8k/Gpt
+Mistral7b-Base
+MCSE 0.650 0.777 0.738 0.676
+MCNSE 0.656 0.798 0.745 0.602
+Semantic Entropy 0.661 0.808 0.757 0.692
+CEDegMat 0.693 0.845 0.760 0.625
+SAR 0.695 0.849 0.771 0.643
+Semantic Density 0.523 0.854 0.696 0.633
+MSP 0.684 0.836 0.795 0.663
+$\text{CoCoA}_{MSP}$ 0.713 0.853 0.790 0.720
+PPL 0.657 0.843 0.795 0.592
+$\text{CoCoA}_{PPL}$ 0.697 0.858 0.790 0.615
+MTE 0.627 0.838 0.788 0.606
+$\text{CoCoA}_{MTE}$ 0.689 0.858 0.783 0.632
+Llama8b-Base
+MCSE 0.632 0.763 0.617 0.653
+MCNSE 0.633 0.783 0.629 0.629
+Semantic Entropy 0.653 0.800 0.650 0.669
+CEDegMat 0.716 0.836 0.703 0.633
+SAR 0.698 0.831 0.715 0.659
+Semantic Density 0.534 0.849 0.655 0.666
+MSP 0.647 0.801 0.787 0.627
+$\text{CoCoA}_{MSP}$ 0.706 0.836 0.771 0.686
+PPL 0.648 0.796 0.776 0.635
+$\text{CoCoA}_{PPL}$ 0.702 0.835 0.765 0.658
+MTE 0.624 0.794 0.719 0.644
+$\text{CoCoA}_{MTE}$ 0.701 0.842 0.736 0.673
+Falcon7b-Base
+MCSE 0.662 0.797 0.770 0.589
+MCNSE 0.682 0.826 0.779 0.571
+Semantic Entropy 0.684 0.821 0.788 0.593
+CEDegMat 0.726 0.863 0.785 0.580
+SAR 0.727 0.862 0.806 0.611
+Semantic Density 0.525 0.856 0.734 0.684
+MSP 0.706 0.850 0.820 0.555
+$\text{CoCoA}_{MSP}$ 0.738 0.869 0.815 0.667
+PPL 0.677 0.844 0.820 0.657
+$\text{CoCoA}_{PPL}$ 0.719 0.866 0.815 0.657
+MTE 0.654 0.825 0.818 0.603
+$\text{CoCoA}_{MTE}$ 0.717 0.867 0.810 0.628
+Table 17: AUROC for all models on QA tasks. Main model response obtained by selecting the most probable candidate among stochastically sampled responses
+Appendix H Computational Budget
+Total available computational resources used to produce results in this paper amounted to 12 compute nodes each having 4xA100 40Gb GPUs.
+Total computational budget spent to produce results was around 400 GPU-days.
+Each individual combination of model and dataset amounted roughly to 16 GPU-days on average to obtain all model outputs and hidden states needed for computing the results.